### v2

# Efficient Response Generation Strategy Selection for Fine-Tuning Large Language Models Through Self-Aligned Perplexity

Abstract
Fine-tuning large language models (LLMs) typically relies on producing large sets of input-output pairs. Yet for a given question, there can be many valid outputs. In practice, these outputs are often derived by distilling knowledge from teacher models, and they can vary depending on the specific teacher model or prompting strategy employed. Recent findings show that how these training outputs are generated can significantly affect the performance of the fine-tuned model, raising an important question: how do we pick the best data generation method from among numerous possibilities? Rather than exhaustively training and evaluating on each candidate, this paper proposes a scalable approximate method that assesses a small subset of generated data to estimate its suitability for a specific target LLM. Our central idea is that effective outputs should be familiar to the target LLM. While previous work measures familiarity with perplexity, we find that perplexity might be suboptimal in characterizing “familiarity” through theoretical analysis and practical observations. To address this, we introduce self-aligned perplexity, a novel metric capturing how closely candidate outputs adhere to the target LLM’s own style and reasoning patterns. In this way, we can identify the most effective generation strategy on a small sample, then apply it to produce the complete training set. We demonstrate that training on data generated by the chosen method yields significant improvements across diverse reasoning-focused benchmarks.
1 Introduction
When instruction-tuning an LLM, training data consists of question-response pairs, where multiple valid responses can be generated for the same input. Previous studies (Ren et al., 2024) show that datasets with identical input questions but different responses can lead to varied learning outcomes, even when responses contain similar levels of detail. This raises a key question: how can we construct responses that are most effective for the target LLM?
Prior research has explored improving responses by adding details or rationales, such as structuring ground truth step by step (Hsieh et al., 2023; Ranaldi & Freitas, 2024), incorporating rationales, or enriching responses with additional information (Zhang et al., 2024; Kang et al., 2023; Li et al., 2022). However, recent studies (Ren et al., 2024; Yang et al., 2024) suggest that more details or converting responses to step by step style do not always improve performance and that alignment with the LLM’s linguistic style is crucial.
From our experiment, we observe that, no single response generation strategy works universally across tasks. Thus, we need to creating a method to find out the most effective way to generate response for each task, rather than use a single method for all of the tasks.
Some works (Xu et al., 2024; Kim et al., 2024) attempt to predict the effectiveness of response generation methods by evaluating the entire training dataset. They generate full training datasets using each method and then estimate training effectiveness based on scores computed via algorithms or reward models. However, these approaches are computationally expensive and not scalable.
However, can we predict the effectiveness of each data generation methods efficiently? We observe an interesting phenomenon that each response generation method produces responses with a consistent style, meaning that a small subset of generated examples can effectively represent the entire dataset. Based on this assumption, we propose an efficient ranking pipeline that evaluates a limited number of samples (e.g., 10) to assess the performance of each response generation strategy. This approach uses an alignment estimation function to assign scores to each strategy, enabling us to identify the best-performing method without the need for a full-dataset evaluation.
Previous research (Ren et al., 2024) used perplexity to measure a model’s familiarity with candidate question-answer pairs, proposing that lower-perplexity responses for the same input tend to yield better training performance. However, we found several cases where perplexity-based filtering was ineffective. For instance, responses structured in a step-by-step or redundant language style often have low perplexity but do not necessarily improve training outcomes.When examining the initial response from the target LLM, we note that on some tasks, the probability of the model producing a step-by-step or redundant response in its initial prediction is very low, even though these responses have low perplexity. These findings suggest that perplexity can be easily ”hacked” by response style. Consequently, traditional perplexity alone is insufficient for selecting the best response generation strategy.
To address this, we propose self-aligned perplexity, a refined metric for measuring a model’s familiarity with target responses. The key idea is that a model is most familiar with the data it generates itself. Leveraging this, we modify perplexity computation by incorporating model-generated responses as in-context examples. Specifically, we first have the model produce an initial response, which is then appended to the question as an in-context example. A prompt enforce the model to pay attention to this example when computing perplexity, thereby altering the probability estimation of the candidate response. If the target response deviates significantly from the model’s own generated response—the one it is most familiar with—the model assigns it a lower probability, increasing its perplexity. Our experiments show that self-aligned perplexity outperforms traditional perplexity in selecting effective data generation strategies.
In our experiments, we observe a strong correlation between the proposed indicator and the ranking of training dataset performance. Furthermore, we construct a pool of answer generation strategies and demonstrate that applying our selection criterion leads to significant performance gains compared to the baselines.
2 Related Works
There has been extensive research into what types of data yield the best training outcomes for large language models (LLMs). Previous studies have identified several factors that positively influence model training, such as adding complexity (Xu et al., 2023), adding details (Zhang et al., 2024; Kang et al., 2023; Li et al., 2022), adding diversity (Luo et al., 2023), augmenting ground-truth answers in a step-by-step manner (Hsieh et al., 2023; Ho et al., 2022; Magister et al., 2023; Fu et al., 2023; Ranaldi & Freitas, 2024), and ensuring correctness (Trinh et al., 2024; Ranaldi & Freitas, 2024). However, in practice, these metrics are challenging to measure for a given dataset, making it difficult to determine the quality of training data based on these criteria. Ren et al. (2024) found that familiarity, measured by perplexity, significantly impacts model training. They argue that language models are more familiar with content generated by other language models, which explains why synthetic data is often more effective for training compared to human-generated data for educating LLMs.
Perplexity has been widely used for different purpose in prior research. Perplexity has been used to select prompts (Gonen et al., 2022), showing that prompts with lower perplexity generally lead to better performance in question-answering tasks. It has also been used for selecting pretraining datasets (De la Rosa et al., 2022), detecting AI-generated content (Xu & Sheng, 2024; Hu et al., 2020), and selecting instruction-tuning data from a database (Mekala et al., 2024). Li et al. (2024) modify the perplexity score and propose “IDF” (Instruction Following Difficulty), which is used to select a small pool of challenging data from the original dataset for efficient training. Researchers hypothesized that higher perplexity reflects more challenging data, which can be beneficial for teaching LLMs new knowledge. Unlike these studies, our focus is on finding the best strategy to generate the target responses (y) for a given x.
Two recent works have explored optimal strategies for generating responses, but their approaches diverge from ours. In (Xu et al., 2024), the authors introduce the Compatibility-Adjusted Reward (CAR) to rank training datasets; however, they limit their experiments to two instruction-following datasets and do not consider multiple meta-prompts. Similarly, (Kim et al., 2024) identifies the key factors influencing training outcomes and employs principal component analysis (PCA) to fit a combination of these factors in order to predict the training performance for the same datasets. In contrast, our work presents several key distinctions. First, unlike the previous work, we evaluate only a small subset of generated data rather than complete datasets, enabling more efficient analysis. Second, while previous studies rely on composite metrics that require hyperparameter tuning—potentially risking overfitting on limited data diversity—we employ a single, robust metric validated across diverse tasks. Third, instead of ranking teacher models using a single meta-prompt, we rank models paired with diverse meta-prompts. Lastly, rather than solely demonstrating ranking effectiveness, our method also enhances overall accuracy.
3 Method
This paper aims to efficiently select the most effective answer generation strategy for fine-tuning a target LLM. In what follows, we first present the problem setup, then detail our proposed self-aligned perplexity metric for scoring the outputs from each candidate strategy.
3.1 Problem Definition
Let be a set of candidate answer-generation strategies, where each strategy produces a response for an input . Our goal is to select the strategy that yields the most effective training data for fine-tuning a target model . Since generating the full dataset via the API for every strategy is costly, we evaluate a small subset of size () to estimate how well each strategy’s outputs align with .
3.2 The familiarity hypothesis
The work in Ren et al. (2024) suggests that if the model is more “familiar” with a given response, then the model can learn better with the given response. In Ren et al. (2024), perplexity, which is equivalent to the likelihood of generating a response with the model, is used to measure this familiarity score. In our study, we argue that perplexity might be sub-optimal to measure the familiarity. We suggest that the familiarity can be more precisely (at least theoretically) measured by the following equation:
where is a semantic similarity measure between and a sample response drawn from the model . In plain language, it quantifies how similar a candidate response is to the range of answers that the model might generate. It is straightforward to demonstrate that when , i.e., when only if is exactly identical to , the function becomes equivalent to the likelihood , and hence equivalent to perplexity. Theoretically, using perplexity as a surrogate to measure familiarity fails to account for the variety of responses that may be semantically equivalent to a candidate response, thereby underestimating the familiarity. In practice, this results in assigning an excessively high perplexity to a good candidate response that the model might actually be familiar with, as evidenced by our empirical study in section 6.1 and examples in Table 1.
| Model | Task | Prompt | Accuracy | Perplexity | Self-algined PPL |
| Mistral7B | GSM8K | Step by Step | 0.627 | 2.09 | 2.26 |
| Claud Answer Directly | 0.592 | 1.96 | 2.45 | ||
| Llama3 | MATH | Step by Step | 0.56 | 1.88 | 1.91 |
| Claud Answer Directly | 0.495 | 1.84 | 2.02 | ||
| Qwen2.5 | API-Bank | Step by Step | 0.41 | 3.02 | 2.98 |
| Claud Answer Directly | 0.461 | 3.14 | 2.90 |
3.3 Self-Aligned Perplexity
To address the issue described above, we introduce an alternative surrogate for measuring familiarity by leveraging the concept of in-context learning. Specifically, we use the model’s initial prediction, , as an in-context example of style and then evaluate the likelihood of a candidate response conditioned on this example in the prompt. This approach encourages the model to assign a higher probability to a response if its style closely aligns with that of the model’s generated response.
Traditional Perplexity.
For a candidate response of length , traditional perplexity is defined as:
Self-Aligned Perplexity.
To address this, we use ’s initial prediction, , as an in-context style example. For each input in a subset :
1.
Generate . 111 may generate incorrect responses, which can sometimes cause to mislead the in-context perplexity calculation. To mitigate this, we employ a filtering mechanism to remove incorrect . According to our experimental results, since is primarily used to guide style, its correctness is less critical. Hence, the correctness filter is optional. For more details, please refer to Appendix A.4 and A.7.
2.
For candidate , include (or when ) in the prompt to provide style guidance.
3.
Compute the perplexity of conditioned on this prompt.
Formally,
This self-aligned measure penalizes responses that deviate from ’s own style. For more details, please read Appendix A.1
Selection Criterion.
For each strategy , we calculate:
and select:
Table 1 shows that self-aligned perplexity more reliably selects the optimal strategy across three tasks (GSM8K, MATH, and API-Bank). For example, on GSM8K, Mistral7B achieves its best accuracy with the step-by-step strategy—correctly favored by self-aligned perplexity but not by standard perplexity.
4 Benchmark Construction
In the section, we shows how do we use different strategies (distinct prompts and teacher LLMs) in generating the high-quality responses with different styles.
4.1 Target LLMs and APIs
We use Mistral-7B-instruct-V2 (Jiang et al., 2023), Llama3-instruct (Dubey & Abhinav Jauhri, 2024) and Qwen-2.5-7B-Instruct(Qwen et al., 2025) as the target language models . In this paper, we refer to Llama3-instruct, Mistral-7B-instruct-V2, and Qwen-2.5-7B-Instruct as Mistral7B, Llama3, and Qwen2.5, respectively. We use GPT-4o, MiniGPT-4o, and Claude 3.5 APIs as teacher models for response generation. Specifically, we use gpt-4o-mini-2024-07-18 and gpt-4o-2024-08-06 (OpenAI, 2023) from OpenAI, and claude-3-5-sonnet-20240620 (Anthropic, 2023) from Anthropic.
4.2 Datasets
We use English reasoning datasets referenced in the technical reports of LLaMA3 (Dubey & Abhinav Jauhri, 2024), Mistral (Jiang et al., 2023), and Qwen-2.5 (Qwen et al., 2025) (the three target models in our experiments). We select datasets with at least 650 examples that can be evaluated via accuracy. If a dataset lacks sufficient training data, we reconstruct it to contain at least 400 training, 50 validation, and 200 testing examples.
For datasets with subcategories (e.g., MATH, MMLU, MMLU_PRO, API_BANK, AGIEVAL), we choose the most challenging subcategory (i.e., with the lowest reported accuracy). For example, we include moral scenarios from MMLU, Professional Law from MMLU_PRO, Level 3 problems from API_BANK, geometry from MATH, and LogicQA from AGIEVAL; we also incorporate the Algebra subcategory from MATH as in Ren et al. (2024).
Following Ren et al. (2024), we train and evaluate the first 1,000 training and testing examples, generating up to 1,000 training examples per data generation strategy.
In total, our datasets include: Mathematics: GSM8K (Cobbe et al., 2021), MATH (Algebra) and MATH (Geometry) (Hendrycks et al., 2021); Commonsense reasoning: PIQA (Bisk et al., 2020), WinoGrande (Sakaguchi et al., 2021), Hellaswag (Zellers et al., 2019), and ECQA (Aggarwal et al., 2021); Reading comprehension: BoolQ (Clark et al., 2019) and SQuAD (Rajpurkar et al., 2016); Aggregated benchmarks: MMLU (Moral Scenarios) (Hendrycks et al., 2020), MMLU_PRO (Professional Law) (Wang et al., 2024), and AGIEval (LogicQA) (Zhong et al., 2023); Coding: MBPP (Austin et al., 2021); Reasoning: DROP (Dua et al., 2019) and ARC-Challenge (Clark et al., 2018); and Tool-using: API-BANK (Lv 3 problems) (Li et al., 2023). For further details, please refer to Table 15 in the Appendix.
4.3 Data Generation Strategies
Given 1,000 samples, we use different strategies to generate target responses. For a fair comparison, we use the same prompts from (Ren et al., 2024) to generate responses, including GPT-4o Answer Directly, Claud Answer Directly, MiniGPT-4o Answer Directly, Step-by-Step and Rewrite Ground Truth. Besides, we design two new prompts named GPT-4o Examples and Human Examples on our own. Please refer to Appendix A.7 for details on each response construction method.
We provide ground truth to the teacher models and allow up to three attempts for data generation. If the first result is incorrect, we regenerate; otherwise, we stop. The same evaluation script used during testing is applied to check correctness.
5 Experiment
In this section, we treat each generation strategy from Section 4.3, along with other response-selection methods mentioned in the related work section, as baselines. We then compare the average training outcomes of our method against these baselines across all tasks.
When training the target model using a particular data generation strategy, we run experiments with random seeds 0, 1, and 2, and report their average performance in Table 9. For our method, we also repeat the process three times, each time selecting a different subset of size from the training data, and then report the average of these three runs. For instance, one run may use the first 50 samples, another the second 50, and so on, and we combine their results to obtain a final average.
5.1 Hyperparameters
We utilize the identical hyperparameter settings as referenced in (Ren et al., 2024). Specifically, for model fine-tuning, a learning rate of 2e-5, a batch size of 32, and a warm-up phase encompassing 10% of the total training iterations are applied. A cosine annealing schedule is implemented for the learning rate, and only the Q and V matrices of the LoRA parameters are fine-tuned with a rank of 8. All models undergo training and evaluation using half-precision arithmetic.
5.2 Our Method vs. Baselines (Each Data Generation Methods)
Figure 1 compares our method against baselines. The Rewrite Ground Truth method is omitted from the Figure 1 and Figure 2 since it applies only to datasets with CoT groundtruth. Other results represent task averages across all datasets, while Rewrite Ground Truth averages a subset of tasks. Despite its exclusion from the figure, all metrics and tables incorporate this method’s results.
As shown in Figure 1, our method outperforms every data generation strategy. On average, it surpasses the best data generation strategy (Claude) by 0.66%. The average performance of all methods is 62.97%, which is 1.22% lower than ours (64.19%).
5.3 Choosing Data Generation Methods Becomes More Important When PGR Increases
When analyzing the detailed performance record (n = 1000) (Table 9 in Appendix), we notice that there is substantial performance variation between methods for some tasks, while for others, the variation is minimal. For example, when training the Qwen model on API-BANK, using training data generated by Claud Answer Directly strategy results in an evaluation accuracy of 45.7%, whereas training with data generated by GPT-4o Examples strategy yields an evaluation accuracy of 33.9%. In contrast, for tasks like GSM8K and MATH Algebra, the performance gap between the best and worst methods is under 2%.
We observe that the performance variation between methods is minimal when training with 1000 samples (data from Table 9 in Appendix) yields only negligible improvement over training with 100 samples (data from Table 13 in Appendix).
To investigate this relationship, we created Figure 3, where the X-axis represents the performance gain ratio (PGR) on a logarithmic scale, and the Y-axis measures the standard diviation in training effectiveness across methods. The PGR is defined as:
This metric quantifies the improvement in accuracy when increasing the training data from 100 to 1000 examples.
While the right plot demonstrates that when the PGR exceeds 2%, the diversity across generation methods increases significantly, the left plot shows a strong positive correlation between and . Notably, when PGR is below 2%, the diversity is nearly 0.
These two subplots show that the variation in training effectiveness increases as PGR increases. The greater the divergence between the training outcomes, the more important it becomes to select the best response generation strategy.
5.3.1 Our Method Performs Better as PGR Increases
We investigate whether our method becomes more effective as the dataset is filtered using increasingly higher PGR thresholds. Figure 2 compares our method with baselines on subsets where the minimum PGR ranges from (all data) to 0.3. As the PGR threshold increases, the performance gap between our method and the average baseline grows from 1.3% to 3.5%, indicating that our method performs especially well on datasets with higher PGR values.
5.3.2 Comparison with Other Data Selection Metrics
Figure 4 shows that our method consistently outperform other methods in terms of accuracy and weighted spearman correlation. As the PGR threshold increases, this gap widens.
6 Ablation Study
6.1 Why Self-Aligned Perplexity Outperforms Traditional Perplexity
| Target Response style | Model | Task | PPL | ICPPL | ICPPL | ICPPL |
| Step by Step(sbs) | Mistral7B | ECQA | 4.476 | 3.695 | 4.85 | 4.329 |
| GPT4 Answer Directly(cad) | 5.551 | 4.116 | 4.768 | 4.456 | ||
| Redundant(r) | 4.944 | 4.334 | 5.615 | 4.326 | ||
| Step by Step(sbs) | Mistral7B | PIQA | 4.290 | 3.816 | 5.968 | 4.028 |
| GPT4 Answer Directly(cad) | 6.277 | 4.053 | 5.962 | 4.250 | ||
| Redundant(r) | 4.547 | 3.919 | 6.724 | 4.027 |
Failure example of perplexity: As shown in Table 2, on the ECQA and PIQA tasks, responses written in a highly redundant style (see Appendix A.7) receive lower perplexity scores—even though the target model is unlikely to generate such responses—while high-quality GPT-4o Answer Directly responses receive higher perplexity.
Here’s a slightly more concise version with one word removed for brevity:
Effectiveness of in-context example: When we incorporate a style-specific in-context example (e.g., a step-by-step response), in-context perplexity adjusts the model’s probability estimation to favor responses that match the provided style. Our method uses the target model’s own prediction as the in-context example, allowing self-aligned perplexity to better reflect the alignment between the model’s style and the response candidates.
| Weighted Method | Model | Accuracy | Spearman’s |
| Ours | Mistral7B | 0.582 | 0.411 |
| Train-then-test lr = 2e-5 | 0.578 | 0.249 | |
| Train-then-test lr = 2e-4 | 0.579 | 0.440 | |
| Ours | Llama3 | 0.650 | 0.344 |
| Train-then-test lr = 2e-5 | 0.626 | 0.257 | |
| Train-then-test lr = 2e-4 | 0.638 | 0.286 | |
| Ours | Qwen2.5 | 0.723 | 0.494 |
| Train-then-test lr = 2e-5 | 0.710 | 0.160 | |
| Train-then-test lr = 2e-4 | 0.726 | 0.345 | |
| Ours | Average | 0.652 | 0.416 |
| Train-then-test lr = 2e-5 | 0.638 | 0.222 | |
| Train-then-test lr = 2e-4 | 0.648 | 0.357 |
| Model | Accuracy | Weighted Spearman’s | ||||||||
| 0–10 | 10–20 | 20–30 | 0–50 | 0–100 | 0–10 | 10–20 | 20–30 | 0–50 | 0–100 | |
| Mistral7B | 0.583 | 0.584 | 0.570 | 0.583 | 0.583 | 0.396 | 0.310 | 0.257 | 0.412 | 0.411 |
| Llama3 | 0.644 | 0.650 | 0.642 | 0.650 | 0.652 | 0.331 | 0.393 | 0.077 | 0.344 | 0.314 |
| Qwen2.5 | 0.714 | 0.711 | 0.718 | 0.723 | 0.726 | 0.363 | 0.472 | 0.327 | 0.494 | 0.401 |
6.2 Our Method vs. Train-Then-Select
One natural—but computationally expensive—approach to selecting the optimal response generation strategy is to adopt a ”Train-Then-Select” procedure. In this approach, we first generate a small dataset (e.g., 100 samples) using each candidate strategy. For each dataset, we train the target model and evaluate its performance. We then rank the strategies based on these results and choose the best-performing one to generate the remainder of the dataset.
When evaluating ”Train-Then-Select”, we train the target model on 100 samples under two settings:
1. Standard Training: A learning rate of 2e-5 for 20 epochs (matching our main setup). The performance accuracies for each strategy under this settings is reported in the Table 13.
2. Intense Training: A learning rate of 2e-4 for 40 epochs. The performance accuracies for each strategy under this settings is reported in the Table 14.
After ranking the strategies using Train-Then-Select, we compare its performance with our method. As shown in Table 3, despite using less data and requiring no training, validation, or testing computations for strategy selection, our method achieves better accuracy and higher Spearman correlation.
6.3 Stability of Our Methods
As shown in Table 4, our method exhibits improved accuracy as the subset size increases. Generally speaking, the performance of our method is consistent. Some of the small subsets might causes degradtion of the accuracy(highlighted in red), thus we recommend to choose a relatively larger size of subsets. In the table, ”0–10”, ”10–20”, and ”20–30” indicate that we use the first, second, and third groups of 10 training samples, respectively, to compute the ranking. ”0–50” and ”0–100” represent subsets containing the first 50 and 100 samples. According to the table, when the subset is large (e.g., K = 50 or K = 100), the performance is better in terms of accuracy and Spearman correlation. However, when the subset is small (K = 10), the performance of each subset becomes less stable.
6.4 Groundtruth vs. Synthetic Data
As shown in Table 9 of Appendix, when ground truth is provided in natural language (e.g., GSM8K, MATH, ECQA, MBPP), training on ground truth is less effective than training on synthetic data. This is because LLMs are more familiar with LLM-generated data, as demonstrated by Ren et al. (2024). However, when the ground truth is written as a gold label without a CoT inference process, training on gold label can sometimes outperform training on CoT synthetic data within the same domain. However, as shown in Table 7 of Appendix, training on gold labels harms cross-domain performance more than training on synthetic data. Additionally, in real-life scenarios, training on natural language data is crucial, as users expect to see the rationale behind the final prediction made by LLMs.
7 Conclusions
In this paper, we present a novel and scalable approach for selecting the optimal response generation strategy to train large language models. We introduce a new metric, self-aligned perplexity, which more effectively evaluates the alignment between the target model and the response options compared to traditional perplexity. Our experiments show that as the PGR increases, selecting the best response generation method becomes more crucial. We demonstrate that choosing the optimal generation strategy based on self-aligned perplexity can lead to substantial improvements in model performance, particularly on tasks with higher PGR. We hope our work will inspire researchers who use perplexity as a downstream metric and those seeking to generate the most effective responses from a set of inputs for training large language models.
Appendix A Appendix
A.1 The Details of Strategy Selection Through Self-Aligned Perplexity
Perplexity
A straightforward way to gauge familiarity is via perplexity, which is typically defined as the exponentiation of the average negative log-likelihood per token. That is, for a candidate response consisting of tokens, we define:
In-Context Perplexity
To address this, we propose self-aligned perplexity, which incorporates ’s own initial prediction as an in-context example of style. Concretely:
1.
Generate responses, , where each response is generated from the corresponding input . may generate incorrect responses, which can sometimes cause to mislead the in-context perplexity calculation. To mitigate this, we employ a filtering mechanism to remove incorrect . According to our experimental results, since is primarily used to guide style, its correctness is less critical. Hence, the correctness filter is optional. For more details, please refer to Appendix A.4.
2.
For each candidate response , we assign in-context examples in a cyclic manner: the in-context example for the first input is based on the second input ’s generated response, the in-context example for is based on ’s response, and so on, until the in-context example for the last input is based on ’s response. This way, the model focuses on imitating the inference style rather than directly copying the answers.
3.
Construct a prompt, , containing both and a directive such as:
“Here is an example of how to reason and respond. Please strictly follow this style. If you drift, you must correct yourself mid-inference.”
The detailed prompt can be found in Appendix A.7.
4.
Compute perplexity for a candidate with the model forced to follow the self-demonstrated style in or .
Formally, define:
In this way, if is fluent but diverges from ’s reasoning style, the perplexity will be higher. This makes self-aligned perplexity a more reliable measure of style mismatch compared to plain token-level perplexity.
Selection Criterion.
For each strategy , we obtain from the small subset and average:
We then select:
A.2 Can we get performance gain if we simply put all of the response variants together?
| Method | Model | DROP | Hellaswag | API-Bank |
| Best | Mistral7B | 0.743 | 0.675 | 0.559 |
| Avg | 0.726 | 0.646 | 0.446 | |
| Total | 0.740 | 0.738 | 0.555 | |
| Mixture of good | 0.770 | 0.731 | 0.555 | |
| Mixture of good | 0.744 | 0.686 | 0.535 | |
| Average of all | 0.711 | 0.686 | 0.433 | |
| Best | Llama3 | 0.805 | 0.718 | 0.547 |
| Avg | 0.778 | 0.711 | 0.392 | |
| Total | 0.810 | 0.738 | 0.490 | |
| Mixture of good | 0.812 | 0.745 | 0.527 | |
| Mixture of good | 0.804 | 0.728 | 0.490 | |
| Average of all | 0.771 | 0.705 | 0.457 | |
| Best | Qwen2.5 | 0.814 | 0.739 | 0.461 |
| Avg | 0.804 | 0.719 | 0.413 | |
| Total | 0.798 | 0.748 | 0.584 | |
| Mixture of good | 0.824 | 0.738 | 0.584 | |
| Mixture of good | 0.818 | 0.742 | 0.490 | |
| Average of all | 0.778 | 0.712 | 0.412 |
Selecting the optimal data generation strategy remains essential, even when resources or funding are unlimited. As shown in Table 5, simply combining six types of synthetic data (Total ) does not guarantee a performance gain over selecting the best synthetic training data. For example, after training the Llama3 model on API-Bank using all six types of synthetic data, the evaluation accuracy is only 49%, much lower than when selecting the Claude Answer Directly data (54.7%). Indeed, according to Table 5, if we combine the mixture of the top three data generation strategies (Mixture of good ), the performance is almost always better than if we simply combine all of the data together (Total ). This underscores the importance of selecting data generation strategies, even if we can afford large-scale synthetic data generation and training.
A.3 Spearman’s Rank Correlation Coefficient
Spearman’s rank correlation coefficient (Spearman’s ) is a non-parametric measure of rank correlation. Unlike Pearson’s correlation, which assesses the strength of a linear relationship between two continuous variables, Spearman’s focuses on the monotonic ordering of items. This makes it robust to outliers and suitable for evaluating ranking tasks where exact values matter less than the relative order in which items appear.
Formally, given two sets of ranks and for the same set of items, Spearman’s is defined as:
Here, and represent the rank of the -th item in each respective ranking. The numerator captures the sum of squared rank differences. If approaches 1, the two rankings are highly positively correlated; if nears 0, there is no monotonic relationship; and if approaches -1, the two rankings are inversely related.
In our experiments, we use the gold-standard orderings provided by the datasets as and the predicted rankings generated by our model and the train-then-test baseline as . By comparing these rankings via Spearman’s , we directly evaluate how faithfully each method captures the underlying ordering of items, providing a valuable complement to other accuracy-based metrics.
A.4 The Effectiveness of the Correctness Filter for Initial Predictions from the Target LLM
Before computing the semantic similarity between the initial prediction of the target LLM and the response from the teacher model, we apply a correctness filter(made by Qwen2.5. For more details, you can read the Appendix A.7) to remove incorrect predictions from the target LLM. In other words, we only compare the semantic similarity between correct initial predictions from the target LLM and responses from the teacher model. For details about how we make the
How effective is the correctness filter?
According to Table 6, the correctness filter helps to improve both accuracy and weighted spearman correlation.
| Weighted Method | Model | Accuracy | Weighted Spearman’s |
| Ours with filter | Mistral7B | 0.582 | 0.411 |
| Ours without filter | 0.574 | 0.364 | |
| Ours with filter | Llama3 | 0.650 | 0.344 |
| Ours without filter | 0.651 | 0.303 | |
| Ours with filter | Qwen2.5 | 0.723 | 0.494 |
| Ours without filter | 0.714 | 0.432 |
A.5 The Impact of Accuracy of the Synthetic Data on Training Outcomes
In our experiment, we aim to ensure the correctness of generated answers by validating them against ground truth answers. Our research seeks to identify the best strategy for generating the optimal version of an answer. In other words, we can adjust data generation strategies to ensure correctness.
In our experiments, we use ground truth answers to guide the generated answers for nearly all datasets, with the only exceptions being mathematical problems. This follows the setting of the paper to maintain consistency with previous work (Ren et al., 2024). This approach might be acceptable since closed-source APIs tend to generate accurate answers. For GSM8K and Math Algebra, GPT-4o, Claude, and MiniGPT-4o achieve accuracies of 90% or above.
To evaluate the impact of accuracy on training outcomes, we conducted the following experiment. As shown in Table 8, we tested three approaches: training on the full dataset, using only correct predictions, and replacing incorrect predictions with rewritten ground truth. These approaches showed less than a 2% improvement overall. Note that in this experiment, GPT-4 refers to the gpt-4-1106-preview API, rather than the gpt-4o-2024-08-06 API, which was used in all other experiments in the paper. The mathematical capabilities of GPT-4o, GPT-4-Mini, and Claude are similar on Math Algebra tasks. Therefore, we used the gpt-4-1106-preview API, which has a weaker ability to solve Math Algebra problems. The benifit of using it is that it makes more mistakes on GSM8K so that we can better evaluate the influence of accuracy. We used this API once to generate the data and train the model from there.
According to the table, the overall benefit of replacing incorrect examples with rewritten ground truth or removing incorrect examples has minimal impact on the overall training outcomes.
A.6 Data Selection Rationale for the Benchmark
The datasets included in our benchmark, drawn from the Mistral, Llama, and Qwen benchmarks, were selected according to a specific set of rules designed to ensure relevance and suitability. These rules are as follows:
1. Sufficient Dataset Size: We only included datasets where the combined size of the training, validation, and testing sets exceeded 650 samples. This threshold was chosen to ensure sufficient data for robust model evaluation.
2. Accuracy as Evaluation Metric: A key requirement was that the dataset could be evaluated using accuracy as the primary metric. This allows for a clear and quantifiable assessment of model performance.
3. English Question-Answering Format: All selected datasets are in an English question-and-answer format to maintain consistency and focus on English language reasoning abilities.
4. Focus on Reasoning Tasks: The underlying task presented by each dataset must involve reasoning skills. This ensures that the benchmark effectively assesses the models’ ability to reason and infer.
A detailed justification for the inclusion or exclusion of each dataset can be found in Table 15.
A.7 Correctness Filter
Without supervised fine-tuning (SFT), may generate incorrect responses, making cosine similarity calculations between and unreliable. To alleviate this, we introduce a filtering mechanism to filter out the incorrect . We notice that for mathematical problems, the correct final answer typically appears as the last number in . Therefore, for Math-related tasks, we use regular expressions (regex) to extract the last number from the prediction and compare it directly with the ground truth. For other types of problem, we use the Qwen2.5-Instruct 7b model to extract the predicted label from the model output. We then compare this extracted label with the true gold label; if they match, we consider the prediction correct by default. The code we used to extract labels are detailed in Appendix A.8
Response Construction Details
Ground Truth: This strategy uses the original ground-truth responses from the datasets as target outputs. Since our focus is on selecting effective chain-of-thought (CoT) target responses, we apply this method to datasets that include human-annotated CoT reasoning steps, such as GSM8K, MATH, ECQA, MBPP. When human-annotated CoT is unavailable, we use the gold label as ground truth.
GPT-4o Answer Directly, Claud Answer Directly, and MiniGPT-4o Answer Directly generate responses based on questions and the ground truth using GPT-4o, Claude 3.5 and Mini-GPT4, respectively. Rewrite Ground Truth: Direct GPT-4o to restyle the ground truth in its own language. This method is only applicable to GSM8K, MATH Algebra, ECQA. The other tasks’s ground truth consists of target labels without any human-annotated chain-of-thought (CoT) reasoning, making rewriting infeasible. Step-by-Step: instructs GPT-4o to generate step-by-step responses based on questions and ground truth. GPT-4o Examples: To facilitate problem-solving, we provide GPT-4o with two high-quality, expert-selected in-context examples of its own responses. GPT-4o is then tasked with generating new responses based on these examples. Human Examples: To aid GPT-4o in understanding problem-solving for these datasets, we provide two carefully chosen human-written examples as context. GPT-4o then uses these examples to generate new responses. We put more details in Section A.7.1 in Appendix.
Redundant Prompt
We construct redundant prompts to demonstrate that the perplexity of the redundant target responses is lower than that of GPT-4’s answers. Perplexity primarily reflects how fluent the language is and how well the language style aligns with the model, but it places less emphasis on semantic meaning.
Declaration of Independence
This is the part of Declaration of Independence that we use in the experiment in Table 4.
Self-Aligned In-Context Prompt for Perplexity Calculation This prompt shows how we add self - generated initial predictions from other questions as in - context examples for perplexity calculation.
A.7.1 Data Generation Strategies
We instruct GPT-4, Claude 3.5, Mini-GPT4 to generate different of target responses using different target reponse generation strategies.
GPT-4/Claude 3.5/Mini-GPT4 Answer Directly: This prompt is from (Ren et al., 2024). For tasks involving mathematics and coding, we submit the problems from our training dataset directly to GPT-4 or Claude 3.5 to obtain their solutions. In the case of classification tasks, we provide these models with the input questions alongside the correct labels (excluding any human-generated explanations) and utilize their outputs. These generated answers are then paired with the original questions to form the GPT-4/Claude 3.5 Direct Answer Training Dataset.
To ensure that the models develop their own problem-solving and analytical capabilities, we deliberately exclude any solutions or rationales related to math, coding, or classification tasks. This approach prevents the models from simply mimicking the ground truth processes, which could otherwise result in some of GPT-4’s predictions lacking its unique reasoning style. Such mimicry would undermine the reliability of our perplexity measurements, which are designed to evaluate how effectively a language model handles outputs from other models.
The prompt below is designed to guide GPT-4/Claude 3.5 in generating responses without relying on the ground truth solutions:
Rewrite Ground Truth: This prompt is from (Ren et al., 2024). In this approach, we provide GPT-4 and Claude 3.5 with the ground truth data, which includes human-annotated rationales and detailed problem-solving steps. The goal is to have GPT-4 and Claude 3.5 rephrase the ground truth content using their own linguistic styles.
The subsequent prompt guides GPT-4 and Claude 3.5 to generate the GPT-4/Claude 3.5 Response (Rewrite GT) output.
Step-by-step: This prompt is from (Ren et al., 2024). We instruct GPT-4 and Claude 3.5 to methodically address each problem by breaking it down into sequential steps. For tasks involving mathematics and coding, we present the problems directly from our training dataset to these models to obtain their solutions. In classification tasks, we provide GPT-4 and Claude 3.5 with the correct labels (excluding any human-generated explanations) along with the input questions, and then utilize their detailed, step-by-step responses. These generated answers are subsequently paired with the original questions to form the GPT-4/Claude 3.5 Step-by-Step Response (No GT) Dataset.
To ensure that the models develop their own unique problem-solving and analytical approaches, we intentionally exclude the solutions or rationales for the mathematics, coding, or classification tasks. This prevents the models from simply mimicking the problem-solving and analytical methods found in the ground truth data. Including such processes could result in some of GPT-4’s and Claude 3.5’s outputs not reflecting their inherent reasoning styles, thereby compromising the accuracy of our perplexity measurements. These measurements are designed to assess how effectively a language model can handle outputs generated by other language models.
The following prompt directs GPT-4 and Claude 3.5 to generate the GPT-4/Claude 3.5 Step-by-Step Response (No GT) responses.
GPT-4o with GPT-4o Examples: We developed this prompt specifically for the API-Bank and Plan-Bench datasets. This prompt utilizes GPT-4’s own accurate generations as examples to help GPT-4 not only better understand the task but also demonstrate how to solve the problems effectively. The prompt below is an example that we used to generate target responses for the API-Bank dataset.
GPT-4 with Human Written Examples: We developed this prompt specifically for the API-Bank and Plan-Bench datasets. This prompt utilizes human written examples to help GPT-4 not only better understand the task but also demonstrate how to solve the problems effectively. The prompt below is an example that we used to generate target responses for the API-Bank dataset.
A.8 Prompt Used to Extract Labels from Predictions
The code below shows how we use Qwen-2.5-Instruct to extract the predicted labels from the predictions.
A.9 AI writing assistant
We use GPT-4o for writing assistant.
| Method | Model Type | training task | GSM8K | Math Algebra | ECQA | SQUAD | DROP | WINOGRANDE |
| Gold Label | Mistral | ECQA | 0.383 | 0.181 | 0.722 | 0.251 | 0.084 | 0.562 |
| GPT-4o Answer Directly | 0.484 | 0.218 | 0.707 | 0.175 | 0.016 | 0.638 | ||
| Gold Label | Mistral | SQUAD | 0.082 | 0.0931 | 0.633 | 0.74 | 0.208 | 0.566 |
| GPT-4o Answer Directly | 0.512 | 0.234 | 0.594 | 0.748 | 0.268 | 0.628 | ||
| Gold Label | Mistral | DROP | 0.076 | 0.097 | 0.621 | 0.561 | 0.628 | 0.578 |
| GPT-4o Answer Directly | 0.542 | 0.241 | 0.602 | 0.546 | 0.736 | 0.638 | ||
| Gold Label | Mistral | WINOGRANDE | 0.381 | 0.172 | 0.625 | 0.166 | 0.042 | 0.742 |
| GPT-4o Answer Directly | 0.477 | 0.219 | 0.569 | 0.106 | 0.016 | 0.713 | ||
| Gold Label | LLAMA3 | ECQA | 0.798 | 0.416 | 0.734 | 0.193 | 0.1 | 0.637 |
| GPT-4o Answer Directly | 0.778 | 0.469 | 0.723 | 0.389 | 0.284 | 0.638 | ||
| Gold Label | LLAMA3 | SQUAD | 0.584 | 0.366 | 0.712 | 0.758 | 0.49 | 0.639 |
| GPT-4o Answer Directly | 0.791 | 0.457 | 0.726 | 0.759 | 0.368 | 0.651 | ||
| Gold Label | LLAMA3 | DROP | 0.144 | 0.169 | 0.674 | 0.574 | 0.738 | 0.582 |
| GPT-4o Answer Directly | 0.776 | 0.507 | 0.703 | 0.555 | 0.786 | 0.626 | ||
| Gold Label | LLAMA3 | WINOGRANDE | 0.776 | 0.445 | 0.717 | 0.226 | 0.162 | 0.766 |
| GPT-4o Answer Directly | 0.775 | 0.485 | 0.721 | 0.305 | 0.238 | 0.695 | ||
| Gold Label | Qwen | ECQA | 0.914 | 0.903 | 0.814 | 0.662 | 0.008 | 0.675 |
| GPT-4o Answer Directly | 0.903 | 0.888 | 0.793 | 0.668 | 0.016 | 0.716 | ||
| Gold Label | Qwen | SQUAD | 0.899 | 0.892 | 0.784 | 0.768 | 0.056 | 0.693 |
| GPT-4o Answer Directly | 0.896 | 0.911 | 0.789 | 0.756 | 0.074 | 0.712 | ||
| Gold Label | Qwen | DROP | 0.788 | 0.904 | 0.799 | 0.701 | 0.664 | 0.711 |
| GPT-4o Answer Directly | 0.911 | 0.903 | 0.792 | 0.741 | 0.806 | 0.701 | ||
| Gold Label | Qwen | WINOGRANDE | 0.893 | 0.904 | 0.78 | 0.651 | 0.004 | 0.725 |
| GPT-4o Answer Directly | 0.902 | 0.896 | 0.798 | 0.68 | 0.022 | 0.721 |
| Dataset | Method | Accuracy and N train | Mistral | Llama3-8B-Chat |
| MATH Algebra | GPT4 preview | 82.5%, 1000 | 0.301 | 0.504 |
| GPT4 only correct | 100%, 825 | 0.293 | 0.501 | |
| GPT4 only correct + rewritten ground truth | 100%, 1000 | 0.293 | 0.500 | |
| MATH Algebra | Claude | 90.1%, 1000 | 0.265 | 0.508 |
| Claude only correct | 100%, 901 | 0.277 | 0.487 | |
| Claude only correct + rewritten ground truth | 100%, 1000 | 0.286 | 0.492 | |
| MATH Algebra | Mini GPT4 | 91.6% , 1000 | 0.313 | 0.523 |
| Mini GPT4 only correct | 100%, 916 | 0.311 | o.523 | |
| Mini GPT4 only correct + rewritten ground truth | 100%, 1000 | 0.326 | 0.539 | |
| GSM8K | GPT4 preview | 92.1%, 1000 | 0.597 | 0.799 |
| GPT4 only correct | 100%, 921 | 0.587 | 0.791 | |
| GPT4 only correct + rewritten ground truth | 100%, 1000 | 0.588 | 0.808 | |
| GSM8K | Claude | 95.6%, 1000 | 0.578 | 0.796 |
| Claude only correct | 100%, 956 | 0.580 | 0.797 | |
| Claude only correct + rewritten ground truth | 100%, 1000 | 0.588 | 0.798 | |
| GSM8K | Mini GPT4 | 89.8% , 1000 | 0.623 | 0.795 |
| Mini GPT4 only correct | 100%, 898 | 0.606 | 0.793 | |
| Mini GPT4 only correct + rewritten ground truth | 100%, 1000 | 0.607 | 0.790 |
| Data Generation Strategy | Model Type | gsm8k | math algebra | math geometry | ecqa | boolq | winogrande | piqa | agieval | squad | arc challenge | drop | mbpp | api bank | hellaswag | mmlu pro law | mmlu moral scenarios |
| gold label | mistral | 0.717 | 0.997 | 0.736 | 0.854 | 0.44 | 0.741 | 0.747 | 0.645 | 0.452 | 0.772 | 0.263 | 0.679 | ||||
| groundtruth | 0.442 | 0.194 | 0.125 | 0.684 | 0.325 | ||||||||||||
| gpt4 | 0.62 | 0.324 | 0.146 | 0.703 | 0.87 | 0.717 | 0.864 | 0.41 | 0.732 | 0.631 | 0.723 | 0.362 | 0.515 | 0.659 | 0.238 | 0.691 | |
| claude | 0.582 | 0.278 | 0.136 | 0.735 | 0.885 | 0.724 | 0.848 | 0.445 | 0.736 | 0.753 | 0.729 | 0.379 | 0.579 | 0.553 | 0.248 | 0.751 | |
| mini gpt4 | 0.619 | 0.306 | 0.151 | 0.708 | 0.882 | 0.695 | 0.868 | 0.427 | 0.732 | 0.772 | 0.735 | 0.348 | 0.43 | 0.663 | 0.205 | 0.659 | |
| step by step | 0.626 | 0.314 | 0.137 | 0.706 | 0.874 | 0.693 | 0.862 | 0.445 | 0.749 | 0.71 | 0.696 | 0.333 | 0.377 | 0.644 | 0.249 | 0.714 | |
| openai human written examples | 0.621 | 0.303 | 0.163 | 0.708 | 0.891 | 0.721 | 0.859 | 0.413 | 0.76 | 0.692 | 0.741 | 0.345 | 0.411 | 0.674 | 0.233 | 0.71 | |
| gpt4 style in context examples | 0.61 | 0.254 | 0.158 | 0.726 | 0.884 | 0.727 | 0.868 | 0.44 | 0.761 | 0.697 | 0.735 | 0.378 | 0.416 | 0.672 | 0.225 | 0.728 | |
| rewrite groundtruth in own words | 0.502 | 0.238 | 0.127 | 0.703 | 0.306 | ||||||||||||
| gold label | llama 3 instruct | 0.737 | 0.979 | 0.761 | 0.852 | 0.432 | 0.756 | 0.766 | 0.742 | 0.507 | 0.772 | 0.332 | 0.639 | ||||
| groundtruth | 0.678 | 0.404 | 0.239 | 0.701 | 0.445 | ||||||||||||
| gpt4 | 0.816 | 0.559 | 0.301 | 0.74 | 0.87 | 0.697 | 0.866 | 0.448 | 0.759 | 0.806 | 0.793 | 0.482 | 0.477 | 0.712 | 0.247 | 0.659 | |
| claude | 0.803 | 0.5 | 0.254 | 0.756 | 0.865 | 0.72 | 0.86 | 0.445 | 0.765 | 0.801 | 0.757 | 0.471 | 0.547 | 0.709 | 0.259 | 0.737 | |
| mini gpt4 | 0.805 | 0.551 | 0.28 | 0.721 | 0.864 | 0.677 | 0.868 | 0.437 | 0.747 | 0.816 | 0.783 | 0.491 | 0.384 | 0.719 | 0.225 | 0.645 | |
| step by step | 0.797 | 0.562 | 0.26 | 0.731 | 0.869 | 0.72 | 0.853 | 0.433 | 0.779 | 0.792 | 0.78 | 0.455 | 0.227 | 0.71 | 0.242 | 0.684 | |
| openai human written examples | 0.81 | 0.547 | 0.283 | 0.735 | 0.893 | 0.717 | 0.867 | 0.44 | 0.766 | 0.804 | 0.807 | 0.477 | 0.347 | 0.706 | 0.229 | 0.667 | |
| gpt4 style in context examples | 0.796 | 0.494 | 0.285 | 0.736 | 0.885 | 0.719 | 0.87 | 0.447 | 0.752 | 0.811 | 0.794 | 0.47 | 0.368 | 0.721 | 0.259 | 0.681 | |
| rewrite groundtruth in own words | 0.742 | 0.444 | 0.241 | 0.727 | 0.437 | ||||||||||||
| gold label | qwen | 0.816 | 0.892 | 0.732 | 0.867 | 0.48 | 0.77 | 0.855 | 0.663 | 0.515 | 0.74 | 0.303 | 0.605 | ||||
| groundtruth | 0.899 | 0.894 | 0.667 | 0.793 | 0.59 | ||||||||||||
| gpt4 | 0.897 | 0.916 | 0.679 | 0.794 | 0.858 | 0.709 | 0.878 | 0.552 | 0.76 | 0.886 | 0.798 | 0.591 | 0.436 | 0.722 | 0.3 | 0.656 | |
| claude | 0.895 | 0.904 | 0.648 | 0.788 | 0.862 | 0.72 | 0.88 | 0.553 | 0.766 | 0.874 | 0.793 | 0.607 | 0.462 | 0.72 | 0.309 | 0.66 | |
| mini gpt4 | 0.904 | 0.904 | 0.654 | 0.787 | 0.87 | 0.712 | 0.882 | 0.555 | 0.763 | 0.891 | 0.821 | 0.642 | 0.379 | 0.701 | 0.308 | 0.664 | |
| step by step | 0.899 | 0.907 | 0.642 | 0.792 | 0.859 | 0.716 | 0.88 | 0.548 | 0.767 | 0.881 | 0.806 | 0.623 | 0.417 | 0.713 | 0.287 | 0.662 | |
| openai human written examples | 0.905 | 0.909 | 0.647 | 0.787 | 0.871 | 0.697 | 0.883 | 0.547 | 0.794 | 0.883 | 0.82 | 0.628 | 0.458 | 0.724 | 0.283 | 0.608 | |
| gpt4 style in context examples | 0.899 | 0.903 | 0.657 | 0.803 | 0.88 | 0.731 | 0.878 | 0.57 | 0.785 | 0.868 | 0.809 | 0.631 | 0.309 | 0.742 | 0.3 | 0.642 | |
| rewrite groundtruth in own words | 0.902 | 0.904 | 0.654 | 0.787 | 0.589 |
| Data Generation Strategy | Model Type | gsm8k | math algebra | math geometry | ecqa | boolq | winogrande | piqa | agieval | squad | arc challenge | drop | mbpp | api bank | hellaswag | mmlu pro law | mmlu moral scenarios |
| gold label | mistral | 0.722 | 0.996 | 0.742 | 0.852 | 0.440 | 0.748 | 0.759 | 0.628 | 0.465 | 0.771 | 0.252 | 0.650 | ||||
| groundtruth | 0.440 | 0.201 | 0.110 | 0.672 | 0.370 | ||||||||||||
| gpt4 | 0.625 | 0.319 | 0.177 | 0.700 | 0.867 | 0.713 | 0.869 | 0.400 | 0.732 | 0.611 | 0.746 | 0.347 | 0.510 | 0.654 | 0.229 | 0.713 | |
| claude | 0.583 | 0.279 | 0.160 | 0.720 | 0.886 | 0.709 | 0.849 | 0.425 | 0.728 | 0.732 | 0.726 | 0.403 | 0.584 | 0.549 | 0.219 | 0.760 | |
| mini gpt4 | 0.627 | 0.291 | 0.148 | 0.710 | 0.873 | 0.688 | 0.877 | 0.420 | 0.740 | 0.775 | 0.726 | 0.363 | 0.433 | 0.663 | 0.183 | 0.643 | |
| step by step | 0.639 | 0.323 | 0.127 | 0.705 | 0.885 | 0.687 | 0.861 | 0.445 | 0.752 | 0.708 | 0.676 | 0.340 | 0.478 | 0.639 | 0.196 | 0.723 | |
| openai human written examples | 0.604 | 0.306 | 0.160 | 0.709 | 0.897 | 0.718 | 0.869 | 0.420 | 0.756 | 0.685 | 0.742 | 0.350 | 0.400 | 0.664 | 0.196 | 0.717 | |
| gpt4 style in context examples | 0.619 | 0.231 | 0.169 | 0.725 | 0.887 | 0.732 | 0.879 | 0.430 | 0.764 | 0.678 | 0.732 | 0.373 | 0.433 | 0.687 | 0.223 | 0.710 | |
| rewrite groundtruth in own words | 0.511 | 0.231 | 0.127 | 0.709 | 0.323 | ||||||||||||
| gold label | llama 3 instruct | 0.734 | 0.978 | 0.766 | 0.855 | 0.435 | 0.761 | 0.764 | 0.738 | 0.502 | 0.777 | 0.312 | 0.630 | ||||
| groundtruth | 0.681 | 0.396 | 0.215 | 0.691 | 0.450 | ||||||||||||
| gpt4 | 0.814 | 0.562 | 0.278 | 0.723 | 0.880 | 0.695 | 0.865 | 0.435 | 0.752 | 0.801 | 0.796 | 0.480 | 0.494 | 0.722 | 0.276 | 0.677 | |
| claude | 0.816 | 0.493 | 0.253 | 0.748 | 0.879 | 0.728 | 0.864 | 0.455 | 0.763 | 0.808 | 0.746 | 0.500 | 0.547 | 0.710 | 0.286 | 0.757 | |
| mini gpt4 | 0.795 | 0.557 | 0.278 | 0.725 | 0.867 | 0.702 | 0.863 | 0.450 | 0.739 | 0.826 | 0.730 | 0.500 | 0.384 | 0.703 | 0.223 | 0.670 | |
| step by step | 0.798 | 0.564 | 0.308 | 0.728 | 0.874 | 0.718 | 0.866 | 0.460 | 0.783 | 0.792 | 0.780 | 0.450 | 0.216 | 0.715 | 0.229 | 0.657 | |
| openai human written examples | 0.811 | 0.547 | 0.266 | 0.736 | 0.891 | 0.719 | 0.864 | 0.450 | 0.770 | 0.809 | 0.808 | 0.457 | 0.355 | 0.699 | 0.269 | 0.640 | |
| gpt4 style in context examples | 0.792 | 0.515 | 0.274 | 0.742 | 0.875 | 0.717 | 0.854 | 0.460 | 0.755 | 0.809 | 0.798 | 0.483 | 0.273 | 0.718 | 0.219 | 0.683 | |
| rewrite groundtruth in own words | 0.729 | 0.443 | 0.241 | 0.715 | 0.417 | ||||||||||||
| gold label | qwen | 0.814 | 0.880 | 0.725 | 0.868 | 0.500 | 0.769 | 0.856 | 0.652 | 0.518 | 0.747 | 0.296 | 0.590 | ||||
| groundtruth | 0.906 | 0.898 | 0.675 | 0.784 | 0.610 | ||||||||||||
| gpt4 | 0.889 | 0.916 | 0.658 | 0.793 | 0.865 | 0.721 | 0.879 | 0.545 | 0.762 | 0.890 | 0.794 | 0.607 | 0.433 | 0.706 | 0.276 | 0.633 | |
| claude | 0.884 | 0.906 | 0.662 | 0.796 | 0.873 | 0.716 | 0.885 | 0.550 | 0.767 | 0.867 | 0.798 | 0.600 | 0.457 | 0.717 | 0.322 | 0.667 | |
| mini gpt4 | 0.905 | 0.904 | 0.654 | 0.782 | 0.865 | 0.704 | 0.881 | 0.535 | 0.760 | 0.891 | 0.818 | 0.633 | 0.396 | 0.704 | 0.299 | 0.657 | |
| step by step | 0.899 | 0.908 | 0.624 | 0.795 | 0.846 | 0.703 | 0.874 | 0.545 | 0.752 | 0.882 | 0.766 | 0.630 | 0.412 | 0.717 | 0.276 | 0.653 | |
| openai human written examples | 0.907 | 0.910 | 0.658 | 0.790 | 0.876 | 0.699 | 0.884 | 0.540 | 0.808 | 0.881 | 0.816 | 0.617 | 0.445 | 0.731 | 0.286 | 0.583 | |
| gpt4 style in context examples | 0.896 | 0.902 | 0.654 | 0.799 | 0.883 | 0.734 | 0.871 | 0.540 | 0.782 | 0.863 | 0.800 | 0.607 | 0.339 | 0.742 | 0.302 | 0.653 | |
| rewrite groundtruth in own words | 0.911 | 0.899 | 0.654 | 0.791 | 0.587 |
| Data Generation Strategy | Model Type | gsm8k | math algebra | math geometry | ecqa | boolq | winogrande | piqa | agieval | squad | arc challenge | drop | mbpp | api bank | hellaswag | mmlu pro law | mmlu moral scenarios |
| gold label | mistral | 0.714 | 0.997 | 0.733 | 0.855 | 0.738 | 0.741 | 0.654 | 0.445 | 0.772 | 0.269 | 0.693 | |||||
| groundtruth | 0.443 | 0.191 | 0.131 | 0.690 | 0.303 | ||||||||||||
| gpt4 | 0.617 | 0.327 | 0.148 | 0.704 | 0.872 | 0.719 | 0.861 | 0.415 | 0.732 | 0.641 | 0.712 | 0.370 | 0.518 | 0.662 | 0.243 | 0.680 | |
| claude | 0.581 | 0.277 | 0.143 | 0.742 | 0.885 | 0.731 | 0.847 | 0.455 | 0.740 | 0.764 | 0.730 | 0.367 | 0.576 | 0.555 | 0.262 | 0.747 | |
| mini gpt4 | 0.615 | 0.314 | 0.148 | 0.707 | 0.886 | 0.698 | 0.863 | 0.430 | 0.728 | 0.771 | 0.740 | 0.340 | 0.429 | 0.663 | 0.216 | 0.667 | |
| step by step | 0.619 | 0.309 | 0.131 | 0.707 | 0.868 | 0.696 | 0.862 | 0.445 | 0.748 | 0.711 | 0.706 | 0.330 | 0.327 | 0.646 | 0.276 | 0.710 | |
| openai human written examples | 0.630 | 0.302 | 0.165 | 0.707 | 0.888 | 0.723 | 0.854 | 0.410 | 0.762 | 0.695 | 0.740 | 0.343 | 0.416 | 0.679 | 0.252 | 0.707 | |
| gpt4 style in context examples | 0.605 | 0.265 | 0.152 | 0.726 | 0.882 | 0.724 | 0.862 | 0.445 | 0.760 | 0.706 | 0.736 | 0.380 | 0.408 | 0.665 | 0.226 | 0.737 | |
| rewrite groundtruth in own words | 0.497 | 0.241 | 0.139 | 0.700 | 0.297 | ||||||||||||
| gold label | llama 3 instruct | 0.738 | 0.979 | 0.759 | 0.850 | 0.430 | 0.754 | 0.767 | 0.744 | 0.510 | 0.769 | 0.342 | 0.643 | ||||
| groundtruth | 0.677 | 0.408 | 0.241 | 0.706 | 0.443 | ||||||||||||
| gpt4 | 0.817 | 0.557 | 0.312 | 0.748 | 0.865 | 0.698 | 0.866 | 0.455 | 0.762 | 0.808 | 0.792 | 0.483 | 0.469 | 0.707 | 0.233 | 0.650 | |
| claude | 0.796 | 0.504 | 0.253 | 0.760 | 0.858 | 0.716 | 0.858 | 0.440 | 0.766 | 0.797 | 0.762 | 0.457 | 0.547 | 0.709 | 0.246 | 0.727 | |
| mini gpt4 | 0.810 | 0.548 | 0.274 | 0.719 | 0.863 | 0.664 | 0.871 | 0.430 | 0.751 | 0.811 | 0.810 | 0.487 | 0.384 | 0.727 | 0.226 | 0.633 | |
| step by step | 0.796 | 0.561 | 0.266 | 0.733 | 0.867 | 0.721 | 0.846 | 0.420 | 0.777 | 0.792 | 0.780 | 0.457 | 0.233 | 0.708 | 0.249 | 0.697 | |
| openai human written examples | 0.809 | 0.547 | 0.291 | 0.735 | 0.894 | 0.716 | 0.868 | 0.435 | 0.764 | 0.801 | 0.806 | 0.487 | 0.343 | 0.709 | 0.209 | 0.680 | |
| gpt4 style in context examples | 0.798 | 0.484 | 0.291 | 0.733 | 0.890 | 0.720 | 0.878 | 0.440 | 0.751 | 0.812 | 0.792 | 0.463 | 0.416 | 0.723 | 0.279 | 0.680 | |
| rewrite groundtruth in own words | 0.749 | 0.445 | 0.253 | 0.733 | 0.447 | ||||||||||||
| gold label | qwen | 0.817 | 0.898 | 0.735 | 0.867 | 0.470 | 0.771 | 0.854 | 0.668 | 0.514 | 0.737 | 0.306 | 0.613 | ||||
| groundtruth | 0.896 | 0.892 | 0.658 | 0.798 | 0.580 | ||||||||||||
| gpt4 | 0.901 | 0.904 | 0.692 | 0.794 | 0.855 | 0.703 | 0.878 | 0.555 | 0.759 | 0.884 | 0.800 | 0.583 | 0.437 | 0.730 | 0.312 | 0.667 | |
| claude | 0.901 | 0.903 | 0.654 | 0.784 | 0.857 | 0.722 | 0.878 | 0.555 | 0.765 | 0.877 | 0.790 | 0.610 | 0.465 | 0.721 | 0.302 | 0.657 | |
| mini gpt4 | 0.903 | 0.904 | 0.662 | 0.789 | 0.872 | 0.716 | 0.882 | 0.565 | 0.765 | 0.891 | 0.822 | 0.647 | 0.371 | 0.700 | 0.312 | 0.667 | |
| step by step | 0.899 | 0.907 | 0.646 | 0.790 | 0.866 | 0.723 | 0.883 | 0.550 | 0.775 | 0.881 | 0.826 | 0.620 | 0.420 | 0.711 | 0.292 | 0.667 | |
| openai human written examples | 0.904 | 0.908 | 0.641 | 0.786 | 0.868 | 0.696 | 0.883 | 0.550 | 0.787 | 0.884 | 0.822 | 0.633 | 0.465 | 0.720 | 0.282 | 0.620 | |
| gpt4 style in context examples | 0.900 | 0.903 | 0.658 | 0.805 | 0.878 | 0.730 | 0.882 | 0.585 | 0.787 | 0.870 | 0.814 | 0.643 | 0.294 | 0.742 | 0.299 | 0.637 | |
| rewrite groundtruth in own words | 0.897 | 0.907 | 0.692 | 0.785 | 0.590 |
| Data Generation Strategy | Model Type | gsm8k | math algebra | math geometry | ecqa | boolq | winogrande | piqa | agieval | squad | arc challenge | drop | mbpp | api bank | hellaswag | mmlu pro law | mmlu moral scenarios |
| gold label | mistral | 0.681 | 0.996 | 0.743 | 0.838 | 0.450 | 0.741 | 0.734 | 0.656 | 0.449 | 0.776 | 0.269 | 0.663 | ||||
| groundtruth | 0.441 | 0.211 | 0.101 | 0.679 | 0.350 | ||||||||||||
| gpt4 | 0.617 | 0.315 | 0.169 | 0.708 | 0.868 | 0.700 | 0.870 | 0.415 | 0.739 | 0.661 | 0.720 | 0.343 | 0.482 | 0.641 | 0.276 | 0.703 | |
| claude | 0.612 | 0.277 | 0.148 | 0.742 | 0.883 | 0.716 | 0.856 | 0.410 | 0.744 | 0.743 | 0.726 | 0.367 | 0.445 | 0.570 | 0.246 | 0.713 | |
| mini gpt4 | 0.622 | 0.320 | 0.177 | 0.703 | 0.865 | 0.688 | 0.855 | 0.435 | 0.740 | 0.768 | 0.708 | 0.353 | 0.429 | 0.670 | 0.219 | 0.697 | |
| step by step | 0.622 | 0.322 | 0.139 | 0.709 | 0.866 | 0.697 | 0.843 | 0.430 | 0.763 | 0.700 | 0.714 | 0.360 | 0.298 | 0.661 | 0.219 | 0.720 | |
| openai human written examples | 0.614 | 0.323 | 0.156 | 0.701 | 0.900 | 0.718 | 0.855 | 0.405 | 0.754 | 0.663 | 0.748 | 0.363 | 0.408 | 0.679 | 0.246 | 0.720 | |
| gpt4 style in context examples | 0.606 | 0.251 | 0.165 | 0.712 | 0.884 | 0.724 | 0.860 | 0.420 | 0.771 | 0.711 | 0.748 | 0.373 | 0.449 | 0.673 | 0.266 | 0.737 | |
| rewrite groundtruth in own words | 0.506 | 0.222 | 0.135 | 0.703 | 0.327 | ||||||||||||
| gold label | llama 3 instruct | 0.735 | 0.980 | 0.760 | 0.865 | 0.445 | 0.757 | 0.762 | 0.740 | 0.465 | 0.784 | 0.329 | 0.663 | ||||
| groundtruth | 0.696 | 0.415 | 0.228 | 0.690 | 0.413 | ||||||||||||
| gpt4 | 0.806 | 0.553 | 0.278 | 0.733 | 0.864 | 0.697 | 0.865 | 0.450 | 0.742 | 0.824 | 0.748 | 0.487 | 0.445 | 0.725 | 0.233 | 0.683 | |
| claude | 0.789 | 0.489 | 0.257 | 0.734 | 0.866 | 0.685 | 0.846 | 0.450 | 0.759 | 0.800 | 0.770 | 0.507 | 0.547 | 0.716 | 0.243 | 0.743 | |
| mini gpt4 | 0.795 | 0.536 | 0.287 | 0.733 | 0.866 | 0.690 | 0.869 | 0.450 | 0.754 | 0.796 | 0.686 | 0.467 | 0.367 | 0.709 | 0.246 | 0.640 | |
| step by step | 0.800 | 0.551 | 0.245 | 0.719 | 0.884 | 0.707 | 0.865 | 0.460 | 0.767 | 0.782 | 0.792 | 0.467 | 0.245 | 0.697 | 0.269 | 0.653 | |
| openai human written examples | 0.796 | 0.529 | 0.287 | 0.736 | 0.884 | 0.714 | 0.863 | 0.450 | 0.757 | 0.808 | 0.800 | 0.460 | 0.367 | 0.689 | 0.223 | 0.680 | |
| gpt4 style in context examples | 0.800 | 0.500 | 0.283 | 0.729 | 0.876 | 0.709 | 0.856 | 0.440 | 0.767 | 0.809 | 0.816 | 0.480 | 0.433 | 0.708 | 0.252 | 0.683 | |
| rewrite groundtruth in own words | 0.754 | 0.431 | 0.291 | 0.715 | 0.457 | ||||||||||||
| gold label | qwen | 0.818 | 0.887 | 0.724 | 0.867 | 0.495 | 0.774 | 0.861 | 0.652 | 0.539 | 0.740 | 0.302 | 0.590 | ||||
| groundtruth | 0.901 | 0.910 | 0.675 | 0.758 | 0.623 | ||||||||||||
| gpt4 | 0.897 | 0.892 | 0.654 | 0.791 | 0.858 | 0.710 | 0.883 | 0.540 | 0.777 | 0.882 | 0.788 | 0.603 | 0.433 | 0.703 | 0.306 | 0.607 | |
| claude | 0.881 | 0.916 | 0.641 | 0.785 | 0.859 | 0.735 | 0.877 | 0.540 | 0.762 | 0.872 | 0.798 | 0.597 | 0.461 | 0.732 | 0.292 | 0.690 | |
| mini gpt4 | 0.902 | 0.904 | 0.658 | 0.778 | 0.875 | 0.711 | 0.880 | 0.555 | 0.760 | 0.890 | 0.798 | 0.613 | 0.396 | 0.696 | 0.309 | 0.677 | |
| step by step | 0.886 | 0.907 | 0.679 | 0.770 | 0.859 | 0.715 | 0.869 | 0.560 | 0.767 | 0.867 | 0.792 | 0.597 | 0.404 | 0.711 | 0.332 | 0.677 | |
| openai human written examples | 0.900 | 0.887 | 0.646 | 0.794 | 0.881 | 0.707 | 0.872 | 0.540 | 0.802 | 0.895 | 0.804 | 0.603 | 0.449 | 0.724 | 0.306 | 0.640 | |
| gpt4 style in context examples | 0.911 | 0.908 | 0.650 | 0.792 | 0.875 | 0.728 | 0.891 | 0.535 | 0.790 | 0.866 | 0.824 | 0.577 | 0.318 | 0.734 | 0.316 | 0.643 | |
| rewrite groundtruth in own words | 0.905 | 0.899 | 0.637 | 0.799 | 0.600 |
| Data Generation Strategy | Model Type | gsm8k | math algebra | math geometry | ecqa | boolq | winogrande | piqa | agieval | squad | arc challenge | drop | mbpp | api bank | hellaswag | mmlu pro law | mmlu moral scenarios |
| gold label | mistral | 0.627 | 0.869 | 0.608 | 0.814 | 0.430 | 0.582 | 0.704 | 0.482 | 0.220 | 0.625 | 0.153 | 0.420 | ||||
| groundtruth | 0.420 | 0.205 | 0.101 | 0.591 | 0.267 | ||||||||||||
| gpt4 | 0.513 | 0.231 | 0.101 | 0.596 | 0.837 | 0.636 | 0.790 | 0.345 | 0.333 | 0.624 | 0.244 | 0.317 | 0.249 | 0.269 | 0.166 | 0.380 | |
| claude | 0.505 | 0.215 | 0.110 | 0.634 | 0.837 | 0.627 | 0.804 | 0.400 | 0.290 | 0.630 | 0.250 | 0.340 | 0.257 | 0.284 | 0.179 | 0.413 | |
| mini gpt4 | 0.511 | 0.223 | 0.097 | 0.619 | 0.845 | 0.644 | 0.782 | 0.360 | 0.404 | 0.633 | 0.210 | 0.337 | 0.253 | 0.223 | 0.183 | 0.343 | |
| step by step | 0.494 | 0.247 | 0.080 | 0.593 | 0.845 | 0.636 | 0.765 | 0.355 | 0.314 | 0.618 | 0.092 | 0.317 | 0.265 | 0.254 | 0.183 | 0.403 | |
| openai human written examples | 0.504 | 0.230 | 0.118 | 0.611 | 0.853 | 0.639 | 0.811 | 0.355 | 0.467 | 0.578 | 0.280 | 0.317 | 0.257 | 0.316 | 0.166 | 0.517 | |
| gpt4 style in context examples | 0.500 | 0.245 | 0.114 | 0.560 | 0.845 | 0.649 | 0.789 | 0.340 | 0.312 | 0.611 | 0.124 | 0.337 | 0.208 | 0.295 | 0.183 | 0.423 | |
| rewrite groundtruth in own words | 0.450 | 0.214 | 0.110 | 0.603 | 0.317 | ||||||||||||
| gold label | llama 3 instruct | 0.710 | 0.852 | 0.636 | 0.789 | 0.395 | 0.680 | 0.764 | 0.620 | 0.082 | 0.610 | 0.196 | 0.200 | ||||
| groundtruth | 0.794 | 0.460 | 0.249 | 0.691 | 0.407 | ||||||||||||
| gpt4 | 0.791 | 0.491 | 0.266 | 0.686 | 0.802 | 0.634 | 0.801 | 0.430 | 0.504 | 0.760 | 0.410 | 0.480 | 0.082 | 0.592 | 0.223 | 0.387 | |
| claude | 0.804 | 0.492 | 0.266 | 0.699 | 0.806 | 0.640 | 0.821 | 0.450 | 0.495 | 0.739 | 0.420 | 0.483 | 0.082 | 0.608 | 0.233 | 0.430 | |
| mini gpt4 | 0.797 | 0.477 | 0.257 | 0.710 | 0.800 | 0.621 | 0.823 | 0.425 | 0.509 | 0.751 | 0.400 | 0.497 | 0.086 | 0.589 | 0.229 | 0.373 | |
| step by step | 0.808 | 0.496 | 0.219 | 0.702 | 0.818 | 0.626 | 0.799 | 0.435 | 0.565 | 0.743 | 0.488 | 0.477 | 0.110 | 0.608 | 0.199 | 0.407 | |
| openai human written examples | 0.809 | 0.472 | 0.257 | 0.720 | 0.810 | 0.630 | 0.815 | 0.440 | 0.564 | 0.747 | 0.414 | 0.500 | 0.078 | 0.600 | 0.236 | 0.387 | |
| gpt4 style in context examples | 0.800 | 0.434 | 0.266 | 0.695 | 0.793 | 0.638 | 0.808 | 0.455 | 0.429 | 0.762 | 0.344 | 0.497 | 0.090 | 0.582 | 0.229 | 0.407 | |
| rewrite groundtruth in own words | 0.813 | 0.480 | 0.262 | 0.718 | 0.447 | ||||||||||||
| gold label | qwen | 0.791 | 0.843 | 0.677 | 0.875 | 0.465 | 0.703 | 0.877 | 0.334 | 0.220 | 0.702 | 0.306 | 0.387 | ||||
| groundtruth | 0.913 | 0.918 | 0.679 | 0.792 | 0.603 | ||||||||||||
| gpt4 | 0.908 | 0.898 | 0.692 | 0.802 | 0.831 | 0.711 | 0.863 | 0.560 | 0.661 | 0.891 | 0.092 | 0.637 | 0.237 | 0.697 | 0.326 | 0.580 | |
| claude | 0.911 | 0.912 | 0.679 | 0.788 | 0.837 | 0.718 | 0.876 | 0.550 | 0.652 | 0.894 | 0.114 | 0.640 | 0.237 | 0.691 | 0.282 | 0.577 | |
| mini gpt4 | 0.902 | 0.914 | 0.667 | 0.791 | 0.852 | 0.720 | 0.884 | 0.550 | 0.660 | 0.891 | 0.068 | 0.600 | 0.237 | 0.702 | 0.296 | 0.583 | |
| step by step | 0.909 | 0.919 | 0.599 | 0.802 | 0.848 | 0.708 | 0.870 | 0.545 | 0.682 | 0.879 | 0.062 | 0.617 | 0.224 | 0.690 | 0.309 | 0.563 | |
| openai human written examples | 0.900 | 0.918 | 0.671 | 0.789 | 0.842 | 0.718 | 0.864 | 0.545 | 0.681 | 0.887 | 0.090 | 0.597 | 0.237 | 0.706 | 0.322 | 0.563 | |
| gpt4 style in context examples | 0.918 | 0.914 | 0.679 | 0.798 | 0.836 | 0.710 | 0.865 | 0.535 | 0.678 | 0.888 | 0.050 | 0.627 | 0.196 | 0.696 | 0.326 | 0.567 | |
| rewrite groundtruth in own words | 0.910 | 0.916 | 0.667 | 0.782 | 0.590 |
| Data Generation Strategy | Model Type | gsm8k | math algebra | math geometry | ecqa | boolq | winogrande | piqa | agieval | squad | arc challenge | drop | mbpp | api bank | hellaswag | mmlu pro law | mmlu moral scenarios |
| gold label | mistral | 0.681 | 0.870 | 0.694 | 0.830 | 0.420 | 0.730 | 0.726 | 0.620 | 0.486 | 0.737 | 0.236 | 0.677 | ||||
| groundtruth | 0.409 | 0.186 | 0.093 | 0.638 | 0.293 | ||||||||||||
| gpt4 | 0.586 | 0.270 | 0.152 | 0.672 | 0.864 | 0.686 | 0.821 | 0.455 | 0.649 | 0.736 | 0.670 | 0.340 | 0.404 | 0.623 | 0.233 | 0.687 | |
| claude | 0.554 | 0.237 | 0.122 | 0.663 | 0.858 | 0.701 | 0.855 | 0.405 | 0.690 | 0.760 | 0.662 | 0.360 | 0.400 | 0.619 | 0.243 | 0.710 | |
| mini gpt4 | 0.514 | 0.266 | 0.152 | 0.705 | 0.850 | 0.674 | 0.847 | 0.425 | 0.670 | 0.739 | 0.666 | 0.357 | 0.359 | 0.651 | 0.233 | 0.623 | |
| step by step | 0.575 | 0.235 | 0.131 | 0.662 | 0.853 | 0.667 | 0.842 | 0.415 | 0.691 | 0.746 | 0.646 | 0.327 | 0.286 | 0.575 | 0.233 | 0.593 | |
| openai human written examples | 0.536 | 0.278 | 0.156 | 0.674 | 0.874 | 0.665 | 0.850 | 0.435 | 0.700 | 0.764 | 0.698 | 0.340 | 0.302 | 0.628 | 0.229 | 0.677 | |
| gpt4 style in context examples | 0.548 | 0.222 | 0.156 | 0.658 | 0.879 | 0.681 | 0.864 | 0.430 | 0.676 | 0.741 | 0.650 | 0.333 | 0.343 | 0.628 | 0.203 | 0.687 | |
| rewrite groundtruth in own words | 0.443 | 0.202 | 0.101 | 0.330 | |||||||||||||
| gold label | llama 3 instruct | 0.705 | 0.866 | 0.675 | 0.847 | 0.430 | 0.727 | 0.773 | 0.684 | 0.494 | 0.682 | 0.299 | 0.633 | ||||
| groundtruth | 0.683 | 0.404 | 0.211 | 0.679 | 0.430 | ||||||||||||
| gpt4 | 0.798 | 0.529 | 0.257 | 0.731 | 0.864 | 0.679 | 0.845 | 0.440 | 0.729 | 0.815 | 0.734 | 0.470 | 0.424 | 0.711 | 0.246 | 0.683 | |
| claude | 0.805 | 0.495 | 0.224 | 0.712 | 0.834 | 0.694 | 0.857 | 0.420 | 0.744 | 0.789 | 0.742 | 0.467 | 0.677 | 0.226 | 0.693 | ||
| mini gpt4 | 0.807 | 0.504 | 0.278 | 0.719 | 0.852 | 0.674 | 0.858 | 0.445 | 0.746 | 0.795 | 0.744 | 0.473 | 0.335 | 0.676 | 0.266 | 0.630 | |
| step by step | 0.779 | 0.528 | 0.198 | 0.690 | 0.874 | 0.683 | 0.863 | 0.435 | 0.736 | 0.797 | 0.708 | 0.457 | 0.253 | 0.688 | 0.233 | 0.620 | |
| openai human written examples | 0.772 | 0.483 | 0.249 | 0.712 | 0.873 | 0.678 | 0.853 | 0.405 | 0.726 | 0.789 | 0.772 | 0.473 | 0.302 | 0.674 | 0.262 | 0.640 | |
| gpt4 style in context examples | 0.794 | 0.488 | 0.283 | 0.712 | 0.861 | 0.690 | 0.859 | 0.440 | 0.729 | 0.770 | 0.754 | 0.473 | 0.380 | 0.702 | 0.259 | 0.693 | |
| rewrite groundtruth in own words | 0.693 | 0.415 | 0.232 | 0.430 | |||||||||||||
| gold label | qwen | 0.820 | 0.883 | 0.704 | 0.858 | 0.480 | 0.747 | 0.849 | 0.642 | 0.457 | 0.725 | 0.339 | 0.563 | ||||
| groundtruth | 0.867 | 0.896 | 0.637 | 0.823 | 0.523 | ||||||||||||
| gpt4 | 0.897 | 0.890 | 0.620 | 0.787 | 0.859 | 0.709 | 0.881 | 0.545 | 0.743 | 0.882 | 0.808 | 0.617 | 0.388 | 0.687 | 0.362 | 0.690 | |
| claude | 0.882 | 0.890 | 0.616 | 0.790 | 0.869 | 0.738 | 0.867 | 0.555 | 0.766 | 0.871 | 0.810 | 0.603 | 0.527 | 0.702 | 0.316 | 0.750 | |
| mini gpt4 | 0.889 | 0.912 | 0.624 | 0.794 | 0.867 | 0.719 | 0.887 | 0.530 | 0.750 | 0.891 | 0.772 | 0.580 | 0.429 | 0.707 | 0.289 | 0.640 | |
| step by step | 0.902 | 0.899 | 0.586 | 0.788 | 0.868 | 0.731 | 0.878 | 0.545 | 0.737 | 0.881 | 0.788 | 0.630 | 0.339 | 0.715 | 0.309 | 0.677 | |
| openai human written examples | 0.892 | 0.899 | 0.616 | 0.783 | 0.874 | 0.727 | 0.883 | 0.565 | 0.776 | 0.875 | 0.824 | 0.590 | 0.392 | 0.694 | 0.233 | 0.643 | |
| gpt4 style in context examples | 0.896 | 0.899 | 0.637 | 0.782 | 0.864 | 0.720 | 0.881 | 0.550 | 0.764 | 0.868 | 0.832 | 0.620 | 0.335 | 0.752 | 0.302 | 0.677 | |
| rewrite groundtruth in own words | 0.899 | 0.892 | 0.646 | 0.583 |
| Benchmark Name | Data Name | Chosen/Not Chosen | Why not chosen |
| Mistral 7B | Winogrande | ||
| PIQA | |||
| GSM8K | |||
| MATH | |||
| MBPP | |||
| MMLU | |||
| AGIEVAL | |||
| ARC Challenge | |||
| BoolQ | |||
| Hellaswag | |||
| CommonsenseQA | not a reasoning task | ||
| BBH | In github, it says this dataset can never used in training. | ||
| SIQA | not a reasoning task | ||
| OpenbookQA | not a reasoning task | ||
| ARC Easy | We already choose ARC Challenge | ||
| NaturalQuestions | It evaluates world knowledge instead of reasoning ability | ||
| TriviaQA | It evaluates world knowledge instead of reasoning ability | ||
| QuAC | this is a multiturn, muti context qa dataset. evaluation is too hard | ||
| Llama 3 | MMLU | ||
| MMLU_Pro | |||
| GSM8K | |||
| MATH | |||
| AGIEVAL | |||
| ARC CHALLENGE | |||
| DROP | |||
| API-BANK | |||
| IFEval | less than 650 data | ||
| HumanEval+ | less than 650 data | ||
| BFCL | (subcategory) less than 650 data | ||
| Nexus | Unable to find the dataset | ||
| GPQA | less than 650 data | ||
| HumanEval | less than 650 data | ||
| ZeroSCROLLS/QuALITY | This dataset evaluating model’s long context QA ability. The input is too long thus is hard to train. | ||
| InfiniteBench/En.MC | This dataset evaluating model’s long context QA ability. The input is too long thus is hard to train. | ||
| NIH/Multi-needle | Long context QA task. The input is too long thus is hard to train. Llama already achieves 98.8% accuracy with zero-shot setting. | ||
| Qwen2.5 | MMLU | ||
| MMLU Pro | |||
| MBPP | |||
| ARC CHALLENGE | |||
| GSM8K | |||
| MATH | |||
| WindoGrande | |||
| HellaSwag | |||
| MMLU stem | (subcategory) less than 650 data | ||
| TruthfulQA | not reasoning task | ||
| GPQA | less than 650 data | ||
| TheoremQA | the data set is tooooo challenging for GPT-4o. it does not have the ability to be a teacher for this task. | ||
| HumanEval | less than 650 data | ||
| HumanEval+ | less than 650 data | ||
| MMLU redux | (subcategory) less than 650 data | ||
| BBH | In github, it says this dataset can never used in training. | ||
| MBPP+ | less than 650 data | ||
| MultiPL-E | (subcategory) less than 650 data |

### v3

@@
-Efficient Response Generation Strategy Selection for Fine-Tuning Large Language Models Through Self-Aligned Perplexity
+Efficiently Selecting Response Generation Strategy by Self-Aligned Perplexity for Fine-Tuning LLMs
 Abstract
@@
 Our central idea is that effective outputs should be familiar to the target LLM.
-While previous work measures familiarity with perplexity, we find that perplexity might be suboptimal in characterizing “familiarity” through theoretical analysis and practical observations.
+While previous work measures familiarity with perplexity, we find that perplexity might be suboptimal in characterizing “familiarity” through empirical analyses and practical observations.
 To address this, we introduce self-aligned perplexity, a novel metric capturing how closely candidate outputs adhere to the target LLM’s own style and reasoning patterns.
 In this way, we can identify the most effective generation strategy on a small sample, then apply it to produce the complete training set.
-We demonstrate that training on data generated by the chosen method yields significant improvements across diverse reasoning-focused benchmarks.
+We demonstrate that training on data generated by the chosen method yields significant improvements across diverse reasoning-focused benchmarks, particularly in cases where different candidate methods lead to highly divergent training outcomes.
+Our implementation is publicly available at https://github.com/XuanRen4470/SPPL.
+Efficiently Selecting Response Generation Strategy by Self-Aligned Perplexity for Fine-Tuning LLMs
+Xuan Ren††thanks: Equal contribution.
+University of Adelaide xuan.ren@adelaide.edu.au Qi Chen* University of Adelaide Lingqiao Liu1††thanks: Corresponding author.
+University of Adelaide lingqiao.liu@adelaide.edu.au
 1 Introduction
@@
 This raises a key question: how can we construct responses that are most effective for the target LLM?
-Prior research has explored improving responses by adding details or rationales, such as structuring ground truth step by step (Hsieh et al., 2023; Ranaldi & Freitas, 2024), incorporating rationales, or enriching responses with additional information (Zhang et al., 2024; Kang et al., 2023; Li et al., 2022).
+Prior research has explored improving responses by adding details or rationales, such as structuring ground truth step by step (Hsieh et al., 2023; Ranaldi and Freitas, 2024), incorporating rationales, or enriching responses with additional information (Zhang et al., 2024; Kang et al., 2023; Li et al., 2022).
 However, recent studies (Ren et al., 2024; Yang et al., 2024) suggest that more details or converting responses to step by step style do not always improve performance and that alignment with the LLM’s linguistic style is crucial.
-From our experiment, we observe that, no single response generation strategy works universally across tasks.
-Thus, we need to creating a method to find out the most effective way to generate response for each task, rather than use a single method for all of the tasks.
-Some works (Xu et al., 2024; Kim et al., 2024) attempt to predict the effectiveness of response generation methods by evaluating the entire training dataset.
+In our experiment, we observe that no single response generation strategy works universally across tasks.
+Thus, we need to create a method to find out the most effective way to generate responses for each task, rather than a single method for all tasks.
+The concurrent works (Xu et al., 2024; Kim et al., 2024) attempt to predict the effectiveness of response generation methods by evaluating the entire training dataset.
 They generate full training datasets using each method and then estimate training effectiveness based on scores computed via algorithms or reward models.
@@
 We observe an interesting phenomenon that each response generation method produces responses with a consistent style, meaning that a small subset of generated examples can effectively represent the entire dataset.
-Based on this assumption, we propose an efficient ranking pipeline that evaluates a limited number of samples (e.g., 10) to assess the performance of each response generation strategy.
+Based on this assumption, we propose an efficient ranking pipeline that evaluates a limited number of samples (e.g., 50) to assess the performance of each response generation strategy.
 This approach uses an alignment estimation function to assign scores to each strategy, enabling us to identify the best-performing method without the need for a full-dataset evaluation.
@@
 However, we found several cases where perplexity-based filtering was ineffective.
-For instance, responses structured in a step-by-step or redundant language style often have low perplexity but do not necessarily improve training outcomes.When examining the initial response from the target LLM, we note that on some tasks, the probability of the model producing a step-by-step or redundant response in its initial prediction is very low, even though these responses have low perplexity.
-These findings suggest that perplexity can be easily ”hacked” by response style.
-Consequently, traditional perplexity alone is insufficient for selecting the best response generation strategy.
+For instance, responses structured in a step-by-step or redundant style often exhibit low perplexity but do not necessarily improve training outcomes.
+While some candidate responses (e.g., step-by-step or redundant ones) may achieve low perplexity under the target LLM, the model itself rarely generates such responses when producing answers freely.
+This suggests that low perplexity does not always indicate alignment with the model’s inherent reasoning style.
+These findings suggest that perplexity can be "hacked" by response style.
+Thus, traditional perplexity alone is insufficient for selecting the best response generation strategy.
 To address this, we propose self-aligned perplexity, a refined metric for measuring a model’s familiarity with target responses.
@@
 Leveraging this, we modify perplexity computation by incorporating model-generated responses as in-context examples.
-Specifically, we first have the model produce an initial response, which is then appended to the question as an in-context example.
-A prompt enforce the model to pay attention to this example when computing perplexity, thereby altering the probability estimation of the candidate response.
+Specifically, we first have the model produce initial responses, which is then appended to the question as in-context examples.
+A prompt enforce the model to pay attention to these examples when computing perplexity, thereby altering the probability estimation of the candidate response.
 If the target response deviates significantly from the model’s own generated response—the one it is most familiar with—the model assigns it a lower probability, increasing its perplexity.
@@
 In our experiments, we observe a strong correlation between the proposed indicator and the ranking of training dataset performance.
-Furthermore, we construct a pool of answer generation strategies and demonstrate that applying our selection criterion leads to significant performance gains compared to the baselines.
+Furthermore, we construct a pool of answer generation strategies and demonstrate that applying our selection criterion leads to significant performance gains compared to the baselines—especially in scenarios where different data-generation methods produce highly divergent outcomes
 2 Related Works
 There has been extensive research into what types of data yield the best training outcomes for large language models (LLMs).
-Previous studies have identified several factors that positively influence model training, such as adding complexity (Xu et al., 2023), adding details (Zhang et al., 2024; Kang et al., 2023; Li et al., 2022), adding diversity (Luo et al., 2023), augmenting ground-truth answers in a step-by-step manner (Hsieh et al., 2023; Ho et al., 2022; Magister et al., 2023; Fu et al., 2023; Ranaldi & Freitas, 2024), and ensuring correctness (Trinh et al., 2024; Ranaldi & Freitas, 2024).
+Previous studies have identified several factors that positively influence model training, such as adding complexity (Xu et al., 2023), adding details (Zhang et al., 2024; Kang et al., 2023; Li et al., 2022), adding diversity (Luo et al., 2023), augmenting ground-truth answers in a step-by-step manner (Hsieh et al., 2023; Ho et al., 2022; Magister et al., 2023; Fu et al., 2023; Ranaldi and Freitas, 2024), and ensuring correctness (Trinh et al., 2024; Ranaldi and Freitas, 2024).
 However, in practice, these metrics are challenging to measure for a given dataset, making it difficult to determine the quality of training data based on these criteria.
@@
 (2024) found that familiarity, measured by perplexity, significantly impacts model training.
-They argue that language models are more familiar with content generated by other language models, which explains why synthetic data is often more effective for training compared to human-generated data for educating LLMs.
 Perplexity has been widely used for different purpose in prior research.
 Perplexity has been used to select prompts (Gonen et al., 2022), showing that prompts with lower perplexity generally lead to better performance in question-answering tasks.
-It has also been used for selecting pretraining datasets (De la Rosa et al., 2022), detecting AI-generated content (Xu & Sheng, 2024; Hu et al., 2020), and selecting instruction-tuning data from a database (Mekala et al., 2024).
+It has also been used for selecting pretraining datasets (De la Rosa et al., 2022), detecting AI-generated content (Xu and Sheng, 2024; Hu et al., 2020), and selecting instruction-tuning data from a database (Mekala et al., 2024).
 Li et al.
-(2024) modify the perplexity score and propose “IDF” (Instruction Following Difficulty), which is used to select a small pool of challenging data from the original dataset for efficient training.
-Researchers hypothesized that higher perplexity reflects more challenging data, which can be beneficial for teaching LLMs new knowledge.
-Unlike these studies, our focus is on finding the best strategy to generate the target responses (y) for a given x.
-Two recent works have explored optimal strategies for generating responses, but their approaches diverge from ours.
-In (Xu et al., 2024), the authors introduce the Compatibility-Adjusted Reward (CAR) to rank training datasets; however, they limit their experiments to two instruction-following datasets and do not consider multiple meta-prompts.
-Similarly, (Kim et al., 2024) identifies the key factors influencing training outcomes and employs principal component analysis (PCA) to fit a combination of these factors in order to predict the training performance for the same datasets.
-In contrast, our work presents several key distinctions.
-First, unlike the previous work, we evaluate only a small subset of generated data rather than complete datasets, enabling more efficient analysis.
-Second, while previous studies rely on composite metrics that require hyperparameter tuning—potentially risking overfitting on limited data diversity—we employ a single, robust metric validated across diverse tasks.
-Third, instead of ranking teacher models using a single meta-prompt, we rank models paired with diverse meta-prompts.
-Lastly, rather than solely demonstrating ranking effectiveness, our method also enhances overall accuracy.
+(2024) modify the perplexity score and propose “IFD” (Instruction Following Difficulty), which is used to select a small pool of challenging data from the original dataset for efficient training.
+Researchers hypothesize that higher perplexity indicates more challenging data, which can be beneficial for teaching LLMs new knowledge.
+In addition, perplexity or confidence-based curricula have been explored for NMT (Kocmi and Bojar, 2017) and general text generation (Platanios et al., 2019), where harder (high-perplexity) data are introduced progressively to improve sample efficiency.
+Unlike these studies, our focus is on identifying the best strategy to generate target responses (y) for a given input (x), rather than selecting difficult (x, y) pairs for training language models.
+Recent efforts have begun to ask which teacher model produces the most useful synthetic targets.
+Xu et al.
+(2024) introduce a Compatibility-Adjusted Reward (CAR) and judge its quality by the Spearman correlation between CAR scores and downstream accuracy on two instruction-following datasets, each evaluated with a single meta-prompt.
+Kim et al.
+(2024) study nine datasets spanning mathematics, coding, and general instructions; they correlate several corpus statistics with training gains and combine them with principal-component analysis to rank teacher models.
+Our study differs in four key respects.
+First, we estimate a strategy’s quality from only a small sample of its outputs, making synthetic data generation and evaluation far more affordable.
+Second, we target accuracy improvement, not just rank correlation.
+Third, we experiment on a much broader benchmark: 17 diverse tasks plus 6 Plan-Bench planning tasks.
+Fourth, we evaluate our method on datasets generated using diverse meta-prompts, explicitly accounting for prompt variability.
 3 Method
@@
 Let be a set of candidate answer-generation strategies, where each strategy produces a response for an input .
-Our goal is to select the strategy that yields the most effective training data for fine-tuning a target model .
+Our goal is to select the strategy that yields the most effective training data to fine-tune a target model .
 Since generating the full dataset via the API for every strategy is costly, we evaluate a small subset of size () to estimate how well each strategy’s outputs align with .
@@
 (2024) suggests that if the model is more “familiar” with a given response, then the model can learn better with the given response.
-In Ren et al.
-(2024), perplexity, which is equivalent to the likelihood of generating a response with the model, is used to measure this familiarity score.
-In our study, we argue that perplexity might be sub-optimal to measure the familiarity.
-We suggest that the familiarity can be more precisely (at least theoretically) measured by the following equation:
+In their work, perplexity, which is correlated to the likelihood of generating a response with the model, is used to measure this familiarity score.
+In our study, we argue that perplexity is sub-optimal to measure familiarity.
+We suggest that familiarity can be more precisely measured by this equation:
 where is a semantic similarity measure between and a sample response drawn from the model .
@@
 It is straightforward to demonstrate that when , i.e., when only if is exactly identical to , the function becomes equivalent to the likelihood , and hence equivalent to perplexity.
-Theoretically, using perplexity as a surrogate to measure familiarity fails to account for the variety of responses that may be semantically equivalent to a candidate response, thereby underestimating the familiarity.
-In practice, this results in assigning an excessively high perplexity to a good candidate response that the model might actually be familiar with, as evidenced by our empirical study in section 6.1 and examples in Table 1.
-| Model | Task | Prompt | Accuracy | Perplexity | Self-algined PPL |
-| Mistral7B | GSM8K | Step by Step | 0.627 | 2.09 | 2.26 |
-| Claud Answer Directly | 0.592 | 1.96 | 2.45 | ||
-| Llama3 | MATH | Step by Step | 0.56 | 1.88 | 1.91 |
-| Claud Answer Directly | 0.495 | 1.84 | 2.02 | ||
-| Qwen2.5 | API-Bank | Step by Step | 0.41 | 3.02 | 2.98 |
-| Claud Answer Directly | 0.461 | 3.14 | 2.90 |
+using perplexity as a surrogate to measure familiarity fails to account for the variety of responses that may be semantically equivalent to a candidate response, thereby underestimating the familiarity.
+In practice, this results in assigning an excessively high perplexity to a good candidate response that the model might actually be familiar with, as evidenced by our empirical study in section 6.1.
 3.3 Self-Aligned Perplexity
-To address the issue described above, we introduce an alternative surrogate for measuring familiarity by leveraging the concept of in-context learning.
-Specifically, we use the model’s initial prediction, , as an in-context example of style and then evaluate the likelihood of a candidate response conditioned on this example in the prompt.
-This approach encourages the model to assign a higher probability to a response if its style closely aligns with that of the model’s generated response.
-Traditional Perplexity.
-For a candidate response of length , traditional perplexity is defined as:
-Self-Aligned Perplexity.
-To address this, we use ’s initial prediction, , as an in-context style example. For each input in a subset :
-1.
-Generate .
-111 may generate incorrect responses, which can sometimes cause to mislead the in-context perplexity calculation.
-To mitigate this, we employ a filtering mechanism to remove incorrect .
-According to our experimental results, since is primarily used to guide style, its correctness is less critical.
-Hence, the correctness filter is optional.
-For more details, please refer to Appendix A.4 and A.7.
-2.
-For candidate , include (or when ) in the prompt to provide style guidance.
-3.
-Compute the perplexity of conditioned on this prompt.
-Formally,
-This self-aligned measure penalizes responses that deviate from ’s own style. For more details, please read Appendix A.1
-Selection Criterion.
-For each strategy , we calculate:
-and select:
-Table 1 shows that self-aligned perplexity more reliably selects the optimal strategy across three tasks (GSM8K, MATH, and API-Bank).
-For example, on GSM8K, Mistral7B achieves its best accuracy with the step-by-step strategy—correctly favored by self-aligned perplexity but not by standard perplexity.
+To evaluate the effectiveness of different response generation strategies, we first construct a small calibration set , consisting of the first examples (e.g., ) input questions.
+For each strategy, we generate one candidate response per question .
+Rather than directly measuring the likelihood of a response given a question, as in traditional perplexity, we compute its likelihood under a prompt augmented with in-context exemplars to better reflect stylistic alignment.
+Specifically, for each input , we let the target model generate an initial prediction .
+Since these initial predictions may be incorrect, we evaluate them using the same task-specific script used at test time, and retain only those that are correct.
+These correct model-generated responses form a style pool .
+To measure how well a candidate (generated by a specific strategy) aligns with the model’s own response style, we compute its perplexity under a prompt that includes two in-context exemplars randomly sampled from the style pool (see Appendix 6 for the prompt).
+Each exemplar comes from a different question than () to avoid leaking the answer.
+This measurement corresponds to what we refer to as the self-aligned perplexity (SPPL) of .
+Compared to standard perplexity, which only conditions on the input question, SPPL incorporates stylistic guidance from the model’s own correct outputs.
+Lower SPPL indicates stronger alignment with the model’s preferred generation style.
+For each response generation strategy , we compute the SPPL for all of its generated responses in , and average the scores to obtain a strategy-level value .
+We then rank all strategies by their average SPPL and select the one with the lowest score to generate the full training set.
 4 Benchmark Construction
-In the section, we shows how do we use different strategies (distinct prompts and teacher LLMs) in generating the high-quality responses with different styles.
+In this section, we show how we use different strategies (distinct prompts and teacher LLMs) in generating high-quality responses with different styles.
 4.1 Target LLMs and APIs
-We use Mistral-7B-instruct-V2 (Jiang et al., 2023), Llama3-instruct (Dubey & Abhinav Jauhri, 2024) and Qwen-2.5-7B-Instruct(Qwen et al., 2025) as the target language models .
+We use Mistral-7B-instruct-V2 (Jiang et al., 2023), Llama3-instruct (Dubey and Abhinav Jauhri, 2024) and Qwen-2.5-7B-Instruct(Qwen et al., 2025) as the target language models .
 In this paper, we refer to Llama3-instruct, Mistral-7B-instruct-V2, and Qwen-2.5-7B-Instruct as Mistral7B, Llama3, and Qwen2.5, respectively.
@@
 4.2 Datasets
-We use English reasoning datasets referenced in the technical reports of LLaMA3 (Dubey & Abhinav Jauhri, 2024), Mistral (Jiang et al., 2023), and Qwen-2.5 (Qwen et al., 2025) (the three target models in our experiments).
+We use English reasoning datasets referenced in the technical reports of LLaMA3 (Dubey and Abhinav Jauhri, 2024), Mistral (Jiang et al., 2023), and Qwen-2.5 (Qwen et al., 2025) (the three target models in our experiments).
 We select datasets with at least 650 examples that can be evaluated via accuracy.
 If a dataset lacks sufficient training data, we reconstruct it to contain at least 400 training, 50 validation, and 200 testing examples.
-For datasets with subcategories (e.g., MATH, MMLU, MMLU_PRO, API_BANK, AGIEVAL), we choose the most challenging subcategory (i.e., with the lowest reported accuracy).
+For datasets with subcategories (e.g., MATH, MMLU, MMLU_PRO, API_BANK, AGIEVAL), we choose the challenging subcategory (i.e., with the lowest reported accuracy).
 For example, we include moral scenarios from MMLU, Professional Law from MMLU_PRO, Level 3 problems from API_BANK, geometry from MATH, and LogicQA from AGIEVAL; we also incorporate the Algebra subcategory from MATH as in Ren et al.
@@
 Following Ren et al. (2024), we train and evaluate the first 1,000 training and testing examples, generating up to 1,000 training examples per data generation strategy.
+Main-experiment corpus.
 In total, our datasets include: Mathematics: GSM8K (Cobbe et al., 2021), MATH (Algebra) and MATH (Geometry) (Hendrycks et al., 2021); Commonsense reasoning: PIQA (Bisk et al., 2020), WinoGrande (Sakaguchi et al., 2021), Hellaswag (Zellers et al., 2019), and ECQA (Aggarwal et al., 2021); Reading comprehension: BoolQ (Clark et al., 2019) and SQuAD (Rajpurkar et al., 2016); Aggregated benchmarks: MMLU (Moral Scenarios) (Hendrycks et al., 2020), MMLU_PRO (Professional Law) (Wang et al., 2024), and AGIEval (LogicQA) (Zhong et al., 2023); Coding: MBPP (Austin et al., 2021); Reasoning: DROP (Dua et al., 2019) and ARC-Challenge (Clark et al., 2018); and Tool-using: API-BANK (Lv 3 problems) (Li et al., 2023).
-For further details, please refer to Table 15 in the Appendix.
+More details are in Table 19 (Appendix).
+PlanBench Extension.
+We further evaluate the most challenging subtasks of PlanBench (Valmeekam et al., 2023)—those on which GPT-3 attains an accuracy below 20%.
+The subtasks comprise plan generation, plan optimization, plan verification, plan reuse, plan generalization, and replanning.
+Although we experimented with various prompt formats, Qwen consistently failed to solve any execution problems.
+Since our method relies on generating correct responses for use as in-context examples, we exclude the execution task from our evaluation.
+The remaining six categories remain sufficiently challenging and are not part of our main training benchmarks.
+We use them solely to analyze performance variance when models are trained on tasks that are very challenging.
+How do different response generation strategies affect performance variance under such conditions?
 4.3 Data Generation Strategies
@@
 5 Experiment
-In this section, we treat each generation strategy from Section 4.3, along with other response-selection methods mentioned in the related work section, as baselines.
+| Methods | STD Range | num of recorded data | mistral | llama 3 instruct | qwen | Avg Acc |
+| Upper bound | All Data | 51.0 | 59.39% | 64.44% | 71.41% | 65.08% |
+| Step-by-step | 56.25% | 60.94% | 69.87% | 62.35% | ||
+| GPT-4 ICL examples | 57.29% | 62.03% | 69.90% | 63.07% | ||
+| Human examples | 56.95% | 61.91% | 70.22% | 63.03% | ||
+| Mini-GPT-4 | 56.72% | 61.36% | 70.13% | 62.74% | ||
+| GPT-4 | 56.92% | 62.83% | 69.89% | 63.21% | ||
+| Claude | 57.45% | 62.93% | 70.30% | 63.56% | ||
+| Ours | 58.34% | 63.66% | 70.50% | 64.16% | ||
+| Ours - Claude | All Data | 51.0 | +0.88% | +0.72% | +0.19% | +0.60% |
+| Ours - Avg of Others | +1.41% | +1.66% | +0.45% | +1.17% | ||
+| Ours - Claude | 21.0 | +1.67% | +1.14% | -0.21% | +1.18% | |
+| Ours - Avg of Others | +2.17% | +2.84% | +1.88% | +2.41% | ||
+| Ours - Claude | 10.0 | +3.61% | +1.56% | -1.63% | +2.26% | |
+| Ours - Avg of Others | +3.60% | +6.00% | +3.38% | +4.54% |
+| Methods | STD Range | num of recorded data | mistral | llama 3 instruct | qwen | Avg Acc | Weighted Spearman Pho |
+| Upper bound | All Data | 51.0 | 59.39% | 64.44% | 71.41% | 65.08% | |
+| skywork | 56.32% | 61.64% | 70.26% | 62.74% | 0.271 | ||
+| CAR | 56.36% | 61.80% | 70.31% | 62.82% | 0.279 | ||
+| IFD | 57.23% | 61.65% | 69.85% | 62.91% | 0.191 | ||
+| perplexity | 57.48% | 63.65% | 70.37% | 63.83% | 0.301 | ||
+| Ours | 58.34% | 63.66% | 70.50% | 64.16% | 0.324 | ||
+| Ours - Perplexity | All Data | 51.0 | +0.86% | +0.00% | +0.12% | +0.33% | +0.023 |
+| Ours - Perplexity | STD > 2.00% | 21.0 | +1.62% | +0.08% | +0.50% | +0.80% | +0.032 |
+| Ours - Perplexity | STD > 4.00% | 10.0 | +3.05% | +0.18% | +1.63% | +1.76% | +0.075 |
+In this section, we treat each generation strategy from Section 4.3, and response-selection metrics from the related work section, as baselines.
 We then compare the average training outcomes of our method against these baselines across all tasks.
-When training the target model using a particular data generation strategy, we run experiments with random seeds 0, 1, and 2, and report their average performance in Table 9.
-For our method, we also repeat the process three times, each time selecting a different subset of size from the training data, and then report the average of these three runs.
-For instance, one run may use the first 50 samples, another the second 50, and so on, and we combine their results to obtain a final average.
+There are two benchmark sets, detailed in Section 4.2.
+The first is the Main-experiment corpus, which covers a diverse range of tasks and serves as the primary benchmark for evaluating both the general ranking ability of our metric and the average performance gains achievable by our method.
+Since our goal is efficient data selection, we evaluate each metric using only a small subset of the training data.
+For each method, we repeat the process three times, each time selecting a different subset of size K=50 from the training dataset, and report the average performance across these runs.
+For example, one run may use the first 50 samples, another the second 50, and so on.
+The final result is computed as the average of these three evaluations.
 5.1 Hyperparameters
@@
 All models undergo training and evaluation using half-precision arithmetic.
-5.2 Our Method vs. Baselines (Each Data Generation Methods)
-Figure 1 compares our method against baselines.
-The Rewrite Ground Truth method is omitted from the Figure 1 and Figure 2 since it applies only to datasets with CoT groundtruth.
-Other results represent task averages across all datasets, while Rewrite Ground Truth averages a subset of tasks.
-Despite its exclusion from the figure, all metrics and tables incorporate this method’s results.
-As shown in Figure 1, our method outperforms every data generation strategy.
-On average, it surpasses the best data generation strategy (Claude) by 0.66%.
-The average performance of all methods is 62.97%, which is 1.22% lower than ours (64.19%).
-5.3 Choosing Data Generation Methods Becomes More Important When PGR Increases
-When analyzing the detailed performance record (n = 1000) (Table 9 in Appendix), we notice that there is substantial performance variation between methods for some tasks, while for others, the variation is minimal.
-For example, when training the Qwen model on API-BANK, using training data generated by Claud Answer Directly strategy results in an evaluation accuracy of 45.7%, whereas training with data generated by GPT-4o Examples strategy yields an evaluation accuracy of 33.9%.
-In contrast, for tasks like GSM8K and MATH Algebra, the performance gap between the best and worst methods is under 2%.
-We observe that the performance variation between methods is minimal when training with 1000 samples (data from Table 9 in Appendix) yields only negligible improvement over training with 100 samples (data from Table 13 in Appendix).
-To investigate this relationship, we created Figure 3, where the X-axis represents the performance gain ratio (PGR) on a logarithmic scale, and the Y-axis measures the standard diviation in training effectiveness across methods.
-The PGR is defined as:
-This metric quantifies the improvement in accuracy when increasing the training data from 100 to 1000 examples.
-While the right plot demonstrates that when the PGR exceeds 2%, the diversity across generation methods increases significantly, the left plot shows a strong positive correlation between and .
-Notably, when PGR is below 2%, the diversity is nearly 0.
-These two subplots show that the variation in training effectiveness increases as PGR increases.
-The greater the divergence between the training outcomes, the more important it becomes to select the best response generation strategy.
-5.3.1 Our Method Performs Better as PGR Increases
-We investigate whether our method becomes more effective as the dataset is filtered using increasingly higher PGR thresholds.
-Figure 2 compares our method with baselines on subsets where the minimum PGR ranges from (all data) to 0.3.
-As the PGR threshold increases, the performance gap between our method and the average baseline grows from 1.3% to 3.5%, indicating that our method performs especially well on datasets with higher PGR values.
-5.3.2 Comparison with Other Data Selection Metrics
-Figure 4 shows that our method consistently outperform other methods in terms of accuracy and weighted spearman correlation. As the PGR threshold increases, this gap widens.
+5.2 Evaluation Metrics
+Accuracy.
+For every {model, dataset} pair, we let each ranking metric select the top-ranked response-generation strategy, fine-tune the model on data produced by that strategy, and record the resulting test accuracy.
+We then report the macro average of these accuracies across all evaluated tasks.
+This score answers the practical question: If I trust a metric to choose my training data, how well will my model perform on average?
+Weighted Spearman correlation.
+To measure how closely a metric’s ranking matches the gold ranking, we compute a weighted Spearman coefficient in which each task is weighted by the standard deviation of accuracies obtained from all candidate strategies; tasks whose choice of strategy matters more thus contribute more.
+The exact formula and implementation details are provided in Appendix A.4.
+5.3 Comparison with Baseline Response Generation Strategies
+Table 1 summarizes the average test accuracy obtained when the target model is fine-tuned on data produced by each response-generation strategy.
+For datasets that provide chain-of-thought (CoT) groundtruth, we additionally evaluate the Rewrite Ground Truth strategy.
+As this strategy is only applicable to CoT datasets and some datasets do not have CoT groundtruth, its results are excluded from the table to avoid skewing the overall averages; nevertheless, they are included in every metric that ranks candidate strategies on a per-task basis.
+Effect of task-specific variance.
+Table 13 shows that the performance gap among generation strategies is highly task-dependent: some tasks show differences of several percentage points, while others are nearly insensitive to the chosen strategy.
+To quantify how much our method helps when the choice of generation strategy matters most, we group every {model, dataset} pair by the standard deviation (SD) of accuracies across baselines.
+All tasks include all pairs without filtering.
+High-variance tasks retain only those with .
+Very-high-variance tasks retain only those with .
+In the whole Main-experiment corpus, our approach delivers the highest mean accuracy, exceeding the strongest single baseline (Claude) by 0.60% and the mean of all baselines by 1.17%.
+When we restrict evaluation to high-variance tasks, the average gain of our method over Claude increases to 1.18%; under the very-high-variance filter, this gain further rises to 2.26%.
+Compared with the mean of all baselines, the improvements reach 2.41% and 4.54% on the high-variance and very-high-variance subsets, respectively.
+These results confirm that self-aligned perplexity is particularly valuable in scenarios where candidate generation strategies produce widely divergent training outcomes.
+For a more detailed analysis of performance variance across different tasks and models, please refer to Section 6.2.
+| Methods | STD Range | num of recorded data | mistral | llama 3 instruct | qwen | Avg Acc | Weighted Spearman Pho |
+| Upper bound | All Data | 18.0 | 52.88% | 54.87% | 41.87% | 49.88% | |
+| Step-by-step | 37.56% | 44.48% | 31.06% | 37.70% | |||
+| GPT-4 ICL examples | 40.86% | 45.84% | 36.86% | 41.19% | |||
+| Human examples | 45.01% | 41.82% | 29.89% | 38.91% | |||
+| Mini-GPT-4 | 38.68% | 40.29% | 30.51% | 36.49% | |||
+| GPT-4 | 39.33% | 41.86% | 31.08% | 37.42% | |||
+| Claude | 51.89% | 50.09% | 37.27% | 46.42% | |||
+| Ours | 45.29% | 50.06% | 39.66% | 45.00% | |||
+| Ours - Claude | -6.60% | -0.03% | +2.39% | -1.42% | |||
+| Ours - Avg of Others | +3.07% | +5.99% | +6.89% | +5.31% | |||
+| Upper bound | All Data | 18.0 | 52.88% | 54.87% | 41.87% | 49.88% | |
+| skywork | 43.66% | 46.01% | 36.12% | 41.93% | 0.101 | ||
+| CAR | 41.87% | 44.99% | 35.00% | 40.62% | 0.133 | ||
+| IFD | 52.88% | 49.73% | 38.19% | 46.93% | 0.386 | ||
+| perplexity | 42.69% | 48.50% | 36.44% | 42.54% | 0.198 | ||
+| Ours | 45.29% | 50.06% | 39.66% | 45.00% | 0.241 | ||
+| Ours - Perplexity | +2.59% | +1.56% | +3.22% | +2.46% | +0.043 |
+5.4 Comparison with Alternative Response Selection Metrics
+Table 2 reports results obtained with the same set-up as in Section 5.3, but swapping the ranking metric.
+Across the full Main-experiment corpus, self-aligned perplexity achieves the best mean accuracy and the highest weighted Spearman correlation; standard perplexity is the closest baseline.
+All tasks: Using every training run, our metric surpasses standard perplexity by 0.33% in accuracy and by 0.033 in weighted Spearman .
+High-variance tasks (): The margins widen to 0.80% in accuracy and 0.032 in weighted .
+Very-high-variance tasks (): Gains further increase to 1.76% in accuracy and 0.075 in weighted .
+These results mirror the trend observed in Section 5.3: the larger the performance spread among candidate strategies, the more our metric outperforms conventional perplexity, underscoring its value for selecting high-quality training data.
+5.5 Performance Differences among Response Generation Strategies Can be Very Large
+Candidate response-generation strategies can yield significantly different results depending on the task.
+To illustrate this, we evaluate various strategies on the PlanBench benchmark introduced in Section 4.2, which is designed to be more difficult than standard instruction-following datasets due to its long-horizon, goal-conditioned reasoning requirements.
+As shown in Table 3 and Tabel 20(Appendix), training outcomes vary significantly across methods, underscoring the importance of selecting an appropriate generation strategy.
+Our self-aligned perplexity metric improves accuracy by an average of over standard perplexity and over the mean performance of all strategies.
+The results further demonstrate that, as the optimal model varies across datasets and continues to shift as APIs evolve, model-aware selection metrics like self-aligned perplexity remain critical.
+As shown in Table 3, the average accuracy of IFD is slightly higher than that of our method.
+However, our method shows comparatively better performance on the main tasks, which cover a broader set of 17 tasks (Table 2).
+This suggests that our approach may offer more stable and reliable results when applied across a wider range of tasks.
 6 Ablation Study
 6.1 Why Self-Aligned Perplexity Outperforms Traditional Perplexity
-| Target Response style | Model | Task | PPL | ICPPL | ICPPL | ICPPL |
+| Target Response style | Model | Task | PPL | SPPL | SPPL | SPPL |
 | Step by Step(sbs) | Mistral7B | ECQA | 4.476 | 3.695 | 4.85 | 4.329 |
@@
 | Redundant(r) | 4.547 | 3.919 | 6.724 | 4.027 |
-Failure example of perplexity: As shown in Table 2, on the ECQA and PIQA tasks, responses written in a highly redundant style (see Appendix A.7) receive lower perplexity scores—even though the target model is unlikely to generate such responses—while high-quality GPT-4o Answer Directly responses receive higher perplexity.
-Here’s a slightly more concise version with one word removed for brevity:
-Effectiveness of in-context example: When we incorporate a style-specific in-context example (e.g., a step-by-step response), in-context perplexity adjusts the model’s probability estimation to favor responses that match the provided style.
-Our method uses the target model’s own prediction as the in-context example, allowing self-aligned perplexity to better reflect the alignment between the model’s style and the response candidates.
+Traditional perplexity is sensitive to surface-level stylistic cues, so a low score does not necessarily mean the response “feels” familiar to the model.
+We therefore anchor the metric on the model’s own zero-shot prediction: the closer a candidate lies to this anchor, the more familiar it should be.
+Injecting that prediction as a single in-context example reshapes the probability distribution, yielding a self-aligned perplexity that more faithfully reflects the response’s true familiarity.
+When perplexity fails.
+According to Table 4, On ECQA, a deliberately redundant answer(see Appendix A.6 for how we construct this dataset) scores 4.94 in raw perplexity, while the terser, higher-quality GPT-4-direct answer scores 5.55 (Table 4).
+A similar pattern appears in PIQA (4.55 vs.
+6.28).
+Thus, lower perplexity can sometimes reflect wordiness rather than genuine familiarity with the model’s preferred style.
+How self-aligned perplexity helps.
+According to Table 4, adding a single in-context example can realign the perplexity scores.
+For the GPT-4–style response on ECQA, the raw perplexity (PPL) is 5.551, which is higher than the redundant-style response (4.944).
+After prepending an in-context example drawn from another GPT-4–style answer, the GPT-4 response’s perplexity drops to 4.768.
+In contrast, when the same example is added to the redundant and step-by-step responses, their perplexities increase from 4.944 and 4.476 to 5.615 and 4.850, respectively.
+A similar effect occurs on PIQA: the GPT-4 response on Mistral has an initial perplexity of 6.277, higher than the redundant style (4.570).
+With a GPT-4 in-context example, its perplexity decreases to 5.962, while the redundant style’s perplexity rises to 6.724.
+Across the two tasks, using the model’s own prediction as the in-context anchor consistently lowers the score for its native style by 0.6–1.5 points, restoring the correct ordering and yielding rankings that track downstream fine-tuning gains.
 | Weighted Method | Model | Accuracy | Spearman’s |
-| Ours | Mistral7B | 0.582 | 0.411 |
-| Train-then-test lr = 2e-5 | 0.578 | 0.249 | |
-| Train-then-test lr = 2e-4 | 0.579 | 0.440 | |
-| Ours | Llama3 | 0.650 | 0.344 |
-| Train-then-test lr = 2e-5 | 0.626 | 0.257 | |
-| Train-then-test lr = 2e-4 | 0.638 | 0.286 | |
-| Ours | Qwen2.5 | 0.723 | 0.494 |
-| Train-then-test lr = 2e-5 | 0.710 | 0.160 | |
-| Train-then-test lr = 2e-4 | 0.726 | 0.345 | |
-| Ours | Average | 0.652 | 0.416 |
-| Train-then-test lr = 2e-5 | 0.638 | 0.222 | |
-| Train-then-test lr = 2e-4 | 0.648 | 0.357 |
-| Model | Accuracy | Weighted Spearman’s | ||||||||
-| 0–10 | 10–20 | 20–30 | 0–50 | 0–100 | 0–10 | 10–20 | 20–30 | 0–50 | 0–100 | |
-| Mistral7B | 0.583 | 0.584 | 0.570 | 0.583 | 0.583 | 0.396 | 0.310 | 0.257 | 0.412 | 0.411 |
-| Llama3 | 0.644 | 0.650 | 0.642 | 0.650 | 0.652 | 0.331 | 0.393 | 0.077 | 0.344 | 0.314 |
-| Qwen2.5 | 0.714 | 0.711 | 0.718 | 0.723 | 0.726 | 0.363 | 0.472 | 0.327 | 0.494 | 0.401 |
+| Ours | Mistral7B | 0.583 | 0.317 |
+| TTT (lr=2e-5) | 0.578 | 0.188 | |
+| TTT (lr=2e-4) | 0.579 | 0.557 | |
+| Ours | Llama3 | 0.637 | 0.288 |
+| TTT (lr=2e-5) | 0.613 | 0.168 | |
+| TTT (lr=2e-4) | 0.627 | 0.398 | |
+| Ours | Qwen2.5 | 0.705 | 0.367 |
+| TTT (lr=2e-5) | 0.702 | 0.065 | |
+| TTT (lr=2e-4) | 0.709 | 0.273 | |
+| Ours | Average | 0.642 | 0.324 |
+| TTT (lr=2e-5) | 0.631 | 0.141 | |
+| TTT (lr=2e-4) | 0.638 | 0.409 |
+| 0–10 | 10–20 | 20–30 | 0–30 | 30–60 | 60–90 | 0–50 | 50–100 | 100–150 | |
+| Accuracy (%) | 63.90% | 64.22% | 63.93% | 64.25% | 63.96% | 64.22% | 64.17% | 64.10% | 64.22% |
+| Weighted | 0.289 | 0.363 | 0.290 | 0.327 | 0.286 | 0.269 | 0.317 | 0.288 | 0.367 |
+| 0–100 | 100–200 | 200–300 | 0–200 | 0–300 | - | - | - | - | |
+| Accuracy (%) | 64.18% | 64.09% | 64.08% | 64.13% | 64.13% | - | - | - | - |
+| Weighted | 0.332 | 0.330 | 0.323 | 0.333 | 0.342 | - | - | - | - |
+6.2 Why Do Some Datasets Show Greater Variance in Training Outcomes?
+| Model | gsm8k | math algebra | mmlu | winogrande | piqa | agieval | squad | ecqa | boolq | arc challenge | mmlu pro law | drop | hellaswag | mmlu moral scenarios | math geometry | api bank |
+| qwen | 94.8% / 88.9% | 93.4% / 91.6% | 51.0% / 50.9% | 72.7% / 72.1% | 86.0% / 87.9% | 57.2% / 54.5% | 67.3% / 76.2% | 81.6% / 79.3% | 83.6% / 86.5% | 89.3% / 89.0% | 32.1% / 27.6% | 0.7% / 79.4% | 70.6% / 70.6% | 56.8% / 63.3% | 51.9% / 65.8% | 11.7% / 43.3% |
+| mistral | 39.3% / 62.5% | 15.4% / 31.9% | 37.6% / 42.2% | 44.9% / 71.3% | 69.2% / 86.9% | 33.5% / 40.0% | 8.8% / 73.2% | 57.6% / 70.0% | 82.3% / 86.7% | 73.5% / 61.1% | 19.4% / 22.9% | 0.6% / 74.6% | 39.7% / 65.4% | 37.2% / 71.3% | 5.1% / 17.7% | 1.8% / 51.0% |
+| llama 3 instruct | 85.8% / 81.4% | 58.1% / 56.2% | 43.0% / 48.9% | 68.4% / 69.5% | 81.1% / 86.5% | 43.0% / 43.5% | 24.5% / 75.2% | 72.0% / 72.3% | 78.1% / 88.0% | 79.4% / 80.1% | 25.9% / 27.6% | 15.1% / 79.6% | 56.8% / 72.2% | 33.3% / 67.7% | 18.3% / 27.8% | 0.0% / 49.4% |
+We observed a striking regularity across tasks: whenever enlarging the training set from 100 to 1 000 examples yields little or no accuracy gain, the choice of response-generation method matters equally little.
+Conversely, tasks that continue to improve with more data show pronounced performance gaps between generation strategies.
+Let the improvement ratio be defined as , representing the relative gain from increasing the training size ten-fold.
+Figure 1 plots (x-axis) against the standard deviation of accuracies across generation methods (y-axis).
+A clear positive trend emerges: once the improvement ratio exceeds roughly 2%, the variance among methods rises sharply; below this threshold, it is nearly zero.
+We plot Figure 1 using training results from all tasks in the Main-experiment corpus and PlanBench.
+The results suggest that divergence across generation strategies is greatest exactly when the dataset still offers headroom for improvement.
+On such high-variance tasks, selecting the right response-generation method is critical, underscoring the value of our self-aligned perplexity criterion.
+Figure 1 shows that many red data points lie to the left of the green 2% threshold line, indicating that training with data generated by different response-generation strategies produces negligible variation in training outcomes on Qwen.
+We hypothesize that this low-variance pattern arises because the training set contributes little information beyond Qwen’s existing capabilities.
+This also explains why our method improves accuracy by only 0.45% over the average of all response-generation strategies on Qwen when evaluated on the full dataset.
+In Table 7, we compare zero-shot accuracy with the accuracy of the three models after being trained on GPT-4o direct-answer data for each task.
+As shown, only 43.8% of tasks improve after training, suggesting that Qwen already possesses strong prior knowledge of standard benchmarks for most tasks.
+In contrast, when trained on data with STD > 4%, Qwen achieves a 3.38% accuracy improvement over the average of all other methods.
+On PlanBench, as demonstrated in Section 5.5 and Table 3, our method yields substantial accuracy gains across the six PlanBench subtasks with Qwen: +3.22% over perplexity and +6.89% over the average of all strategies.
+6.3 The effectiveness of Correctness Filter
+Correctness filter is helpful mainly on the tasks that the model is very unfamiliar with.
+Removing the correctness filter has negligible impact on the 17 main tasks, as shown in the Table 9.
+The table shows the accuracy and weighted Spearman correlation deltas between the filtered and unfiltered versions of our method (Ours Ours-w/o-Filter); Overall, skipping the filter—i.e., using the model’s raw predictions—only minimally reduces performance in terms of average accuracy and weighted Spearman correlation.
+However, removing the correctness filter results in a performance drop on PlanBench, as shown in the table 10.
 6.4 Our Method vs. Train-Then-Select
-One natural—but computationally expensive—approach to selecting the optimal response generation strategy is to adopt a ”Train-Then-Select” procedure.
-In this approach, we first generate a small dataset (e.g., 100 samples) using each candidate strategy.
+One natural (but computationally expensive) approach to select the optimal response generation strategy is to adopt a Train-Then-Select (TTS) procedure.
+In this way, we first generate a small dataset (e.g., 100 samples) using each candidate strategy.
 For each dataset, we train the target model and evaluate its performance.
-We then rank the strategies based on these results and choose the best-performing one to generate the remainder of the dataset.
-When evaluating ”Train-Then-Select”, we train the target model on 100 samples under two settings:
-1. Standard Training: A learning rate of 2e-5 for 20 epochs (matching our main setup). The performance accuracies for each strategy under this settings is reported in the Table 13.
-2. Intense Training: A learning rate of 2e-4 for 40 epochs. The performance accuracies for each strategy under this settings is reported in the Table 14.
-After ranking the strategies using Train-Then-Select, we compare its performance with our method.
-As shown in Table 3, despite using less data and requiring no training, validation, or testing computations for strategy selection, our method achieves better accuracy and higher Spearman correlation.
-6.3 Stability of Our Methods
-As shown in Table 4, our method exhibits improved accuracy as the subset size increases.
-Generally speaking, the performance of our method is consistent.
-Some of the small subsets might causes degradtion of the accuracy(highlighted in red), thus we recommend to choose a relatively larger size of subsets.
-In the table, ”0–10”, ”10–20”, and ”20–30” indicate that we use the first, second, and third groups of 10 training samples, respectively, to compute the ranking.
-”0–50” and ”0–100” represent subsets containing the first 50 and 100 samples.
-According to the table, when the subset is large (e.g., K = 50 or K = 100), the performance is better in terms of accuracy and Spearman correlation.
-However, when the subset is small (K = 10), the performance of each subset becomes less stable.
-6.4 Groundtruth vs. Synthetic Data
-As shown in Table 9 of Appendix, when ground truth is provided in natural language (e.g., GSM8K, MATH, ECQA, MBPP), training on ground truth is less effective than training on synthetic data.
+We then rank the strategies based on the results and choose the best-performing one to generate the remainder of the dataset.
+When evaluating TTS, we train the target model on 100 samples under two settings: 1) Standard Training: A learning rate of 2e-5 for 20 epochs (matching our main setup).
+The performance accuracies for each strategy under this setting is in the Table 17.
+2) Intense Training: A learning rate of 2e-4 for 40 epochs.
+The performance accuracies for each strategy under this setting is in Table 18.
+After ranking the strategies using TTS, we compare their performance with ours.
+In Table 5, despite using less data and requiring no training, validation, or testing computations for strategy selection, our method achieves better average accuracy and comparable weighted Spearman correlation.
+6.5 Stability of Our Method
+As shown in Table 6, accuracy generally improves as the subset size grows, and the overall performance is consistent across ranges.
+Small subsets sometimes degrade accuracy (values highlighted in red); thus, we recommend using at least 30 samples or even 50 samples for the best performance.
+Specifically, “0–10”, “10–20”, and “20–30” denote the first, second, and third batches of ten training examples, respectively, while “0–50” and “0–100” correspond to the first 50 and 100 examples.
+When the subset size reaches 50 or more, average accuracy stabilises.
+The weighted Spearman correlation () also increases with larger subsets, but the gains taper off once the subset size exceeds 50.
+7 Conclusions
+In this paper, we present a novel and scalable approach for selecting the optimal response-generation strategy to train large language models.
+We introduce a new metric, self-aligned perplexity, which more effectively evaluates the alignment between a target model and its response options compared to traditional perplexity.
+We demonstrate that choosing the optimal generation strategy based on self-aligned perplexity leads to substantial improvements in model performance, particularly on tasks with high performance variance.
+We hope our work will inspire researchers who use perplexity as a downstream metric or who wish to build the most effective instruction tuning datasets.
+8 Limitations
+While recent open-source “thinking” models(such as DeepSeek-R1-Distill-Qwen-7B) support long chain-of-thought reasoning, our evaluations focus on chat models.
+We used meta prompts to elicit reasoning steps but did not test on “thinking” models.
+We believe our style-alignment approach still applies, though further validation on smaller or differently pretrained models is needed.
+We leave broader scaling studies and extensions to other model families for future work.
+Appendix A Appendix
+A.1 Ground Truth vs. Synthetic Data
+As shown in Table 13 (Appendix), when ground truth is provided in natural language (e.g., GSM8K, MATH, ECQA, MBPP), training on ground truth is less effective than on synthetic data.
 This is because LLMs are more familiar with LLM-generated data, as demonstrated by Ren et al.
 (2024).
-However, when the ground truth is written as a gold label without a CoT inference process, training on gold label can sometimes outperform training on CoT synthetic data within the same domain.
-However, as shown in Table 7 of Appendix, training on gold labels harms cross-domain performance more than training on synthetic data.
-Additionally, in real-life scenarios, training on natural language data is crucial, as users expect to see the rationale behind the final prediction made by LLMs.
-7 Conclusions
-In this paper, we present a novel and scalable approach for selecting the optimal response generation strategy to train large language models.
-We introduce a new metric, self-aligned perplexity, which more effectively evaluates the alignment between the target model and the response options compared to traditional perplexity.
-Our experiments show that as the PGR increases, selecting the best response generation method becomes more crucial.
-We demonstrate that choosing the optimal generation strategy based on self-aligned perplexity can lead to substantial improvements in model performance, particularly on tasks with higher PGR.
-We hope our work will inspire researchers who use perplexity as a downstream metric and those seeking to generate the most effective responses from a set of inputs for training large language models.
-Appendix A Appendix
-A.1 The Details of Strategy Selection Through Self-Aligned Perplexity
-Perplexity
-A straightforward way to gauge familiarity is via perplexity, which is typically defined as the exponentiation of the average negative log-likelihood per token.
-That is, for a candidate response consisting of tokens, we define:
-In-Context Perplexity
-To address this, we propose self-aligned perplexity, which incorporates ’s own initial prediction as an in-context example of style. Concretely:
-1.
-Generate responses, , where each response is generated from the corresponding input .
-may generate incorrect responses, which can sometimes cause to mislead the in-context perplexity calculation.
-To mitigate this, we employ a filtering mechanism to remove incorrect .
-According to our experimental results, since is primarily used to guide style, its correctness is less critical.
-Hence, the correctness filter is optional.
-For more details, please refer to Appendix A.4.
-2.
-For each candidate response , we assign in-context examples in a cyclic manner: the in-context example for the first input is based on the second input ’s generated response, the in-context example for is based on ’s response, and so on, until the in-context example for the last input is based on ’s response.
-This way, the model focuses on imitating the inference style rather than directly copying the answers.
-3.
-Construct a prompt, , containing both and a directive such as:
-“Here is an example of how to reason and respond. Please strictly follow this style. If you drift, you must correct yourself mid-inference.”
-The detailed prompt can be found in Appendix A.7.
-4.
-Compute perplexity for a candidate with the model forced to follow the self-demonstrated style in or .
-Formally, define:
-In this way, if is fluent but diverges from ’s reasoning style, the perplexity will be higher.
-This makes self-aligned perplexity a more reliable measure of style mismatch compared to plain token-level perplexity.
-Selection Criterion.
-For each strategy , we obtain from the small subset and average:
-We then select:
+However, when the ground truth is written as a gold label without a CoT inference process, training on the gold label can sometimes outperform training on CoT synthetic data within the same domain.
+However, in Table 11 (Appendix), training on gold labels harms cross-domain performance more than training on synthetic data.
+Besides, in real-life scenarios, training on natural language data is crucial, as users expect to see the rationale behind the final prediction made by LLMs.
 A.2 Can we get performance gain if we simply put all of the response variants together?
@@
 This underscores the importance of selecting data generation strategies, even if we can afford large-scale synthetic data generation and training.
-A.3 Spearman’s Rank Correlation Coefficient
-Spearman’s rank correlation coefficient (Spearman’s ) is a non-parametric measure of rank correlation.
-Unlike Pearson’s correlation, which assesses the strength of a linear relationship between two continuous variables, Spearman’s focuses on the monotonic ordering of items.
-This makes it robust to outliers and suitable for evaluating ranking tasks where exact values matter less than the relative order in which items appear.
-Formally, given two sets of ranks and for the same set of items, Spearman’s is defined as:
-Here, and represent the rank of the -th item in each respective ranking.
-The numerator captures the sum of squared rank differences.
-If approaches 1, the two rankings are highly positively correlated; if nears 0, there is no monotonic relationship; and if approaches -1, the two rankings are inversely related.
-In our experiments, we use the gold-standard orderings provided by the datasets as and the predicted rankings generated by our model and the train-then-test baseline as .
-By comparing these rankings via Spearman’s , we directly evaluate how faithfully each method captures the underlying ordering of items, providing a valuable complement to other accuracy-based metrics.
-A.4 The Effectiveness of the Correctness Filter for Initial Predictions from the Target LLM
-Before computing the semantic similarity between the initial prediction of the target LLM and the response from the teacher model, we apply a correctness filter(made by Qwen2.5.
-For more details, you can read the Appendix A.7) to remove incorrect predictions from the target LLM.
-In other words, we only compare the semantic similarity between correct initial predictions from the target LLM and responses from the teacher model.
-For details about how we make the
-How effective is the correctness filter?
-According to Table 6, the correctness filter helps to improve both accuracy and weighted spearman correlation.
-| Weighted Method | Model | Accuracy | Weighted Spearman’s |
-| Ours with filter | Mistral7B | 0.582 | 0.411 |
-| Ours without filter | 0.574 | 0.364 | |
-| Ours with filter | Llama3 | 0.650 | 0.344 |
-| Ours without filter | 0.651 | 0.303 | |
-| Ours with filter | Qwen2.5 | 0.723 | 0.494 |
-| Ours without filter | 0.714 | 0.432 |
-A.5 The Impact of Accuracy of the Synthetic Data on Training Outcomes
+A.3 The Impact of Accuracy of the Synthetic Data on Training Outcomes
 In our experiment, we aim to ensure the correctness of generated answers by validating them against ground truth answers.
@@
 According to the table, the overall benefit of replacing incorrect examples with rewritten ground truth or removing incorrect examples has minimal impact on the overall training outcomes.
-A.6 Data Selection Rationale for the Benchmark
+A.4 Weighted Spearman’s Rank Correlation Coefficient
+Spearman’s rank correlation () measures how well two orderings agree, ignoring absolute values.
+Because some of our {model, dataset} pairs exhibit far larger performance gaps among response-generation strategies than others, we assign higher importance to pairs whose choice of strategy matters more.
+We therefore adopt a weighted variant of Spearman’s correlation in which each item is given a non-negative weight .
+Definition.
+Let and be the ranks of the -th item under two orderings and let be its weight. Denote the weighted means
+The weighted Spearman correlation is then the weighted Pearson correlation between the rank vectors:
+Choice of weights.
+For each {model, dataset} pair, we first train the target model on data produced by every candidate response-generation method and record the resulting accuracies.
+The weight is set to the standard deviation of these accuracies.
+Intuitively, tasks in which the strategies yield very different outcomes ( large) are more informative when judging a ranking metric, so they contribute more to .
+Interpretation.
+A value of indicates that the metric produces a ranking almost identical to the gold ranking, with higher-variance tasks influencing the score most strongly.
+Conversely, implies no weighted monotonic relationship, and signals an inverse agreement.
+Throughout the main text and Appendix, all reported “Spearman” results actually correspond to this weighted formulation.
+A.5 Data Selection Rationale for the Benchmark
 The datasets included in our benchmark, drawn from the Mistral, Llama, and Qwen benchmarks, were selected according to a specific set of rules designed to ensure relevance and suitability.
@@
 A detailed justification for the inclusion or exclusion of each dataset can be found in Table 19.
-A.7 Correctness Filter
+A.6 Correctness Filter
 Without supervised fine-tuning (SFT), may generate incorrect responses, making cosine similarity calculations between and unreliable.
@@
 We then compare this extracted label with the true gold label; if they match, we consider the prediction correct by default.
-The code we used to extract labels are detailed in Appendix A.8
+| Methods | STD Range | #Data | Mistral | Llama | Qwen | Acc | Weighted |
+| Ours - Ours w/o Filter | All | 51 | |||||
+| Ours - Ours w/o Filter | 27 | ||||||
+| Ours - Ours w/o Filter | 14 |
+| Methods | STD Range | #Data | Mistral | Llama | Qwen | Acc | Weighted |
+| Ours - Ours w/o Filter | All Data | 18.0 |
 Response Construction Details
@@
 We put more details in Section A.6.2 in Appendix.
+A.6.1 Prompt for Self-Aligned Perplexity
 Redundant Prompt
-We construct redundant prompts to demonstrate that the perplexity of the redundant target responses is lower than that of GPT-4’s answers.
+We construct redundant prompts (shown in the Figure 7) to demonstrate that the perplexity of the redundant target responses is lower than that of GPT-4’s answers.
 Perplexity primarily reflects how fluent the language is and how well the language style aligns with the model, but it places less emphasis on semantic meaning.
-Declaration of Independence
-This is the part of Declaration of Independence that we use in the experiment in Table 4.
-Self-Aligned In-Context Prompt for Perplexity Calculation This prompt shows how we add self - generated initial predictions from other questions as in - context examples for perplexity calculation.
-A.7.1 Data Generation Strategies
+Self-Aligned In-Context Prompt for Perplexity Calculation The prompt shown from the Figure 6 shows how we add self - generated initial predictions from other questions as in - context examples for perplexity calculation.
+A.6.2 Data Generation Strategies
 We instruct GPT-4, Claude 3.5, Mini-GPT4 to generate different of target responses using different target reponse generation strategies.
@@
 Such mimicry would undermine the reliability of our perplexity measurements, which are designed to evaluate how effectively a language model handles outputs from other models.
-The prompt below is designed to guide GPT-4/Claude 3.5 in generating responses without relying on the ground truth solutions:
+The prompt from the Figure 2 below is designed to guide GPT-4/Claude 3.5 in generating responses without relying on the ground truth solutions:
 Rewrite Ground Truth: This prompt is from (Ren et al., 2024).
@@
 The goal is to have GPT-4 and Claude 3.5 rephrase the ground truth content using their own linguistic styles.
-The subsequent prompt guides GPT-4 and Claude 3.5 to generate the GPT-4/Claude 3.5 Response (Rewrite GT) output.
+The subsequent prompt(shown in the Figure 5) guides GPT-4 and Claude 3.5 to generate the GPT-4/Claude 3.5 Response (Rewrite GT) output.
 Step-by-step: This prompt is from (Ren et al., 2024).
@@
 These measurements are designed to assess how effectively a language model can handle outputs generated by other language models.
-The following prompt directs GPT-4 and Claude 3.5 to generate the GPT-4/Claude 3.5 Step-by-Step Response (No GT) responses.
+The following prompt from the Figure 4 directs GPT-4 and Claude 3.5 to generate the GPT-4/Claude 3.5 Step-by-Step Response (No GT) responses.
 GPT-4o with GPT-4o Examples: We developed this prompt specifically for the API-Bank and Plan-Bench datasets.
@@
 The prompt below is an example that we used to generate target responses for the API-Bank dataset.
+The following prompt from the Figure 3 directs GPT-4o to generate responses guided by GPT-4o generated Examples
 GPT-4 with Human Written Examples: We developed this prompt specifically for the API-Bank and Plan-Bench datasets.
@@
 The prompt below is an example that we used to generate target responses for the API-Bank dataset.
-A.8 Prompt Used to Extract Labels from Predictions
-The code below shows how we use Qwen-2.5-Instruct to extract the predicted labels from the predictions.
-A.9 AI writing assistant
-We use GPT-4o for writing assistant.
+The following prompt from the Figure 3 directs GPT-4o to generate responses guided by Human written Examples
+A.7 AI Assistant
+We used GPT-4o as a writing assistant and programming aid for editing purposes.
+A.8 Required Compute Resources
+Each individual training run reported in this paper requires approximately 5–48 GPU hours when using a 40GB A100 GPU.
+We do not recommend you to reproduce every training run, as there are too many experiments.
+Instead, we strongly recommend directly using the reported training outcomes from each table as the final results.
+You can then compute your ranking metrics to evaluate how well your metric aligns with the training outcomes.
+Calculating metrics such as perplexity on a small subset of all of the dataset takes only about 2 hours on a single 40GB A100 GPU.
 | Method | Model Type | training task | GSM8K | Math Algebra | ECQA | SQUAD | DROP | WINOGRANDE |
-| Gold Label | Mistral | ECQA | 0.383 | 0.181 | 0.722 | 0.251 | 0.084 | 0.562 |
-| GPT-4o Answer Directly | 0.484 | 0.218 | 0.707 | 0.175 | 0.016 | 0.638 | ||
-| Gold Label | Mistral | SQUAD | 0.082 | 0.0931 | 0.633 | 0.74 | 0.208 | 0.566 |
-| GPT-4o Answer Directly | 0.512 | 0.234 | 0.594 | 0.748 | 0.268 | 0.628 | ||
-| Gold Label | Mistral | DROP | 0.076 | 0.097 | 0.621 | 0.561 | 0.628 | 0.578 |
-| GPT-4o Answer Directly | 0.542 | 0.241 | 0.602 | 0.546 | 0.736 | 0.638 | ||
-| Gold Label | Mistral | WINOGRANDE | 0.381 | 0.172 | 0.625 | 0.166 | 0.042 | 0.742 |
-| GPT-4o Answer Directly | 0.477 | 0.219 | 0.569 | 0.106 | 0.016 | 0.713 | ||
-| Gold Label | LLAMA3 | ECQA | 0.798 | 0.416 | 0.734 | 0.193 | 0.1 | 0.637 |
-| GPT-4o Answer Directly | 0.778 | 0.469 | 0.723 | 0.389 | 0.284 | 0.638 | ||
-| Gold Label | LLAMA3 | SQUAD | 0.584 | 0.366 | 0.712 | 0.758 | 0.49 | 0.639 |
-| GPT-4o Answer Directly | 0.791 | 0.457 | 0.726 | 0.759 | 0.368 | 0.651 | ||
-| Gold Label | LLAMA3 | DROP | 0.144 | 0.169 | 0.674 | 0.574 | 0.738 | 0.582 |
-| GPT-4o Answer Directly | 0.776 | 0.507 | 0.703 | 0.555 | 0.786 | 0.626 | ||
-| Gold Label | LLAMA3 | WINOGRANDE | 0.776 | 0.445 | 0.717 | 0.226 | 0.162 | 0.766 |
-| GPT-4o Answer Directly | 0.775 | 0.485 | 0.721 | 0.305 | 0.238 | 0.695 | ||
-| Gold Label | Qwen | ECQA | 0.914 | 0.903 | 0.814 | 0.662 | 0.008 | 0.675 |
-| GPT-4o Answer Directly | 0.903 | 0.888 | 0.793 | 0.668 | 0.016 | 0.716 | ||
-| Gold Label | Qwen | SQUAD | 0.899 | 0.892 | 0.784 | 0.768 | 0.056 | 0.693 |
-| GPT-4o Answer Directly | 0.896 | 0.911 | 0.789 | 0.756 | 0.074 | 0.712 | ||
-| Gold Label | Qwen | DROP | 0.788 | 0.904 | 0.799 | 0.701 | 0.664 | 0.711 |
-| GPT-4o Answer Directly | 0.911 | 0.903 | 0.792 | 0.741 | 0.806 | 0.701 | ||
-| Gold Label | Qwen | WINOGRANDE | 0.893 | 0.904 | 0.78 | 0.651 | 0.004 | 0.725 |
-| GPT-4o Answer Directly | 0.902 | 0.896 | 0.798 | 0.68 | 0.022 | 0.721 |
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
 | Dataset | Method | Accuracy and N train | Mistral | Llama3-8B-Chat |
@@
 | MultiPL-E | (subcategory) less than 650 data |
+| Data Generation Strategy | Model Type | plan bench generation | plan bench optimality | plan bench generalization | plan bench replaning | plan bench reuse | plan bench verification |
+| gold label | mistral | 0.607 | 0.448 | 0.807 | 0.815 | 0.945 | 0.318 |
+| groundtruth | |||||||
+| gpt4 | 0.458 | 0.37 | 0.215 | 0.2 | 0.61 | 0.518 | |
+| claude | 0.505 | 0.45 | 0.362 | 0.445 | 0.755 | 0.586 | |
+| mini gpt4 | 0.445 | 0.38 | 0.282 | 0.235 | 0.51 | 0.5 | |
+| step by step | 0.448 | 0.253 | 0.253 | 0.177 | 0.63 | 0.492 | |
+| openai human written examples | 0.485 | 0.43 | 0.302 | 0.28 | 0.73 | 0.474 | |
+| gpt4 style in context examples | 0.37 | 0.232 | 0.43 | 0.207 | 0.665 | 0.56 | |
+| gold label | llama 3 instruct | 0.478 | 0.38 | 0.853 | 0.87 | 0.87 | 0.482 |
+| groundtruth | |||||||
+| gpt4 | 0.463 | 0.36 | 0.265 | 0.225 | 0.72 | 0.49 | |
+| claude | 0.545 | 0.485 | 0.407 | 0.425 | 0.637 | 0.532 | |
+| mini gpt4 | 0.415 | 0.375 | 0.25 | 0.215 | 0.5 | 0.674 | |
+| step by step | 0.475 | 0.39 | 0.287 | 0.232 | 0.732 | 0.564 | |
+| openai human written examples | 0.458 | 0.372 | 0.307 | 0.25 | 0.603 | 0.516 | |
+| gpt4 style in context examples | 0.45 | 0.435 | 0.438 | 0.223 | 0.75 | 0.468 | |
+| gold label | qwen | 0.32 | 0.285 | 0.38 | 0.732 | 0.845 | 0.516 |
+| groundtruth | |||||||
+| gpt4 | 0.237 | 0.253 | 0.145 | 0.2 | 0.515 | 0.538 | |
+| claude | 0.325 | 0.265 | 0.277 | 0.46 | 0.393 | 0.534 | |
+| mini gpt4 | 0.253 | 0.255 | 0.168 | 0.188 | 0.42 | 0.55 | |
+| step by step | 0.265 | 0.182 | 0.135 | 0.17 | 0.577 | 0.536 | |
+| openai human written examples | 0.24 | 0.19 | 0.222 | 0.155 | 0.47 | 0.52 | |
+| gpt4 style in context examples | 0.302 | 0.268 | 0.338 | 0.2 | 0.597 | 0.534 |
+A.9 License of the Dtasets
+All dataset we use are publicly available dataset for research purpose.
+API-BANK (Lv 3 problems) (Li et al., 2023): CC-BY-SA GSM8K (Cobbe et al., 2021): MIT license PIQA (Bisk et al., 2020): unkown BoolQ (Clark et al., 2019):CC BY-SA 3.0 MBPP (Austin et al., 2021):CC BY 4.0 DROP (Dua et al., 2019): CC BY-SA 4.0 ARC-Challenge (Clark et al., 2018):CC BY-SA 4.0 PlanBench (Valmeekam et al., 2023): MIT license MATH (Algebra) and MATH (Geometry) (Hendrycks et al., 2021): MIT license SQuAD (Rajpurkar et al., 2016):SA 4.0 license MMLU(Hendrycks et al., 2020): MIT license WinoGrande (Sakaguchi et al., 2021): Apache-2.0 license Hellaswag (Zellers et al., 2019): MIT license ECQA (Aggarwal et al., 2021): Apache-2.0 license MMLU_PRO (Wang et al., 2024): Apache-2.0 license AGIEval(Zhong et al., 2023): MIT license